### v1

# Can We Catch the Elephant? The Evolvement of Hallucination Evaluation on Natural Language Generation: A Survey

Abstract
Hallucination in Natural Language Generation (NLG) is like the elephant in the room, obvious but often overlooked until recent achievements significantly improved the fluency and grammatical accuracy of generated text. For Large Language Models (LLMs), hallucinations can happen in various downstream tasks and casual conversations, which need accurate assessment to enhance reliability and safety. However, current studies on hallucination evaluation vary greatly, and people still find it difficult to sort out and select the most appropriate evaluation methods. Moreover, as NLP research gradually shifts to the domain of LLMs, it brings new challenges to this direction. This paper provides a comprehensive survey on the evolvement of hallucination evaluation methods, aiming to address three key aspects: 1) Diverse definitions and granularity of facts; 2) The categories of automatic evaluators and their applicability; 3) Unresolved issues and future directions.
1 Introduction
The hallucination problem in Natural Language Generation (NLG) typically refers to situations where the generated text deviates from the facts in source or external knowledge. Like the elephant in the room, this problem has existed since the beginning of the development of NLG. However, in the era of Large Language Models (LLMs), with the rapid improvement of fluency and grammatical accuracy in text generation, this issue has gradually gained more attention. The evaluation of hallucination motivates the development of LLMs to be more reliable and safe. In this paper, we make a comprehensive survey on the hallucination evaluation methods and elaborate them according to the development process, which gives a broad view of current trends and future directions.
The definition of hallucination varies a lot from work to work over time. In the initial stages, hallucination refers to the occurrence of semantically incorrect but grammatically viable content (Lee et al., 2018). This setup is typically rooted in traditional tasks such as Text Summarization and Neural Machine Translation (NMT), which often involve definite source information. The breakthrough came with the advent of LLMs like ChatGPT (OpenAI, 2022), crossing the barriers of conversational systems. These models are capable of processing various text inputs and providing responses. In this case, users have more expectations for the responses, especially the consistency of world facts under open answers. Many previous tasks can also be accomplished within a conversation by prompting LLMs with correct instructions (Ouyang et al., 2022). In the response of LLMs, hallucination occurs now and then, with a mix of faithful and factual problems.
Faithfulness and factuality are two concepts that are closely related when describing hallucinations in the generated text and can be prone to be confusing in some circumstances. In this paper, we add prefixes to both of them for better understanding by introducing Source Faithfulness (SF) and World Factuality (WF)  [^1] In this paper, we also write SF as faithfulness/consistency, and WF as factuality.. SF measures the degree to which the generated output accurately reflects and is consistent with the source input. It ensures that the output stays true to the information provided in the source. SF has a more limited scope, as there are specific source documents that can be used to substantiate and verify the generated text. WF, on the other hand, assesses whether the generated output aligns with general world knowledge and facts. WF is a more expansive and challenging problem as it goes beyond the specific source and considers the broader context of common sense and established knowledge, which are more difficult to comprehensively collect and encode. Recent studies have recognised the critical importance of addressing and improving the SF and WF of generated text.
Whether in existing NLG tasks or the more practical scenarios in the LLMs era, SF and WF problems manifest in different forms. For example, in the NMT task, generated translations detached from the source text are considered factually incorrect (Dale et al., 2023). In the summarization task, generated content usually should belong to the source document to be faithful, but some also argue that certain hallucinations can align with external facts (Dong et al., 2022). In the dialogue task, it is important to balance the hallucination and user experience (Santhanam et al., 2021). In LLM downstream tasks, the issue is often related to the prompt, which is designed for different scenarios and to retrieve the knowledge that existed in LLM. Apart from these, LLM faces unique difficulties, such as updated world information and false-premise questions (Vu et al., 2023; Kasai et al., 2024).
Hallucinations are prevalent across various tasks, as it is not a problem that has been completely defined, yet it intuitively pertains to binary correctness. While humans can readily perceive some hallucinations, the model often struggles to do so. Moreover, in the current development of LLMs regarding reasoning, planning, task adaptability, etc., hallucination evaluation is an issue that requires careful consideration. Previous works have some introductions on benchmarks for LLM hallucinations (Huang et al., 2023; Zhang et al., 2023c; Ji et al., 2023; Huang et al., 2021), but they do not include evaluators before LLMs and more research after, thus resulting in a lack of comparison between developments before and after. Also, they do not specifically focus on evaluators, therefore lacking in-depth analysis of details, as well as their respective strengths and weaknesses.
Our contributions lie in three folds. Firstly, this paper organizes automatic evaluation methods for hallucination in the evolvement timeline, including traditional approaches as well as novel methods employing or assessing LLMs, and conducts analytical comparisons. Secondly, an applicable classification method is proposed, categorizing evaluation methods from a practical perspective based on the scenario. Lastly, recommendations for future hallucination evaluations are provided from multiple perspectives. We also give a brief introduction of related fact-checking and human evaluation works in Appendix A and B.
2 Definition and Taxonomy
To understand how to evaluate hallucination in text generation, we first need to define fact. In previous works, fact granularity is a way to define facts and serves as the foundational basis for assessing hallucination. Fact error classification also can help provide a solid and comprehensive perspective for evaluation.
2.1 Fact Granularity
The definition of fact granularity varies, lacking an agreed-upon labelling scheme (Honovich et al., 2022). Moreover, in generated text, many contents are difficult to have clear definitions. Among previous works, fact granularity can be categorized into the following types.
Token/Word At the token/word level, entities and nouns are the most frequently used granularity, primarily due to their intuitively informative and relative ease of acquisition and comparison. For example, some summarization models tend to forget entities (Maynez et al., 2020; Nan et al., 2021), making up entities that never appeared in the source document.
Span Spans require more annotation than individual tokens or words. Mostly, they are entity-centric, relation-centric, or event-centric, which are closely related to dependency parsing (Goyal & Durrett, 2020; 2021). In evaluators, spans can be used to generate questions (Durmus et al., 2020; Honovich et al., 2021) or check entailment (Kamoi et al., 2023).
Sentence A higher level is sentence, which is coarsely classified whether it contains hallucination or not (Wang et al., 2020; Pagnoni et al., 2021). Sentence-level facts can help obtain an overall perception of the factual content within the sentence.
Passage Passage-level facts can contribute to an overall assessment of the generated text (Fabbri et al., 2021). This encompasses discourse fact connections (Zhang et al., 2023b) and an estimation of passage themes. Moreover, this is crucial for future assessments of the factual accuracy in generating longer texts.
2.2 Fact Error Type
The taxonomies of fact errors vary subjecting to scenarios, some categorize them by the information source as intrinsic and extrinsic hallucinations (Maynez et al., 2020; Goyal & Durrett, 2021; Cao & Wang, 2021), For faithful error types,  Goyal & Durrett (2021) designs entity-, event-, and noun phrase-related errors. Frank (Pagnoni et al., 2021) categorizes with a more fine-grained scale, including PredE, EntE, CircE, CorefE, LinkE, OutE and GramE. We summarize that such errors can be categorized into Source Faithful Error and World Factual Error, in which way it can cover more NLG tasks including traditional ones and LLM scenarios. SF and WF errors are independent as shown in Figure 1.
Figure 1: Source Faithful Error and World Factual Error examples.
3 Evaluation Methods Before LLMs
This section contains the automatic evaluation methods for hallucinations in NLG before LLM, which differ from other evaluators that focus on text similarity or grammar correctness. Specifically, the focus mostly lies on the SF aspect of text generation. While previous works rely on the reference to train the models, the reference itself can suffer a lot of factual errors (Maynez et al., 2020), which hinder the model training from learning faithfulness. Also due to the costly but error-prone reference, many works study reference-free evaluation methods, which only require the input of source documents. To introduce the approaches, we will elaborate from evaluators to benchmarks.
It is noteworthy that a significant portion of these evaluation methods is dedicated to assessing summarization tasks, due to which has a fixed source document as an information reference, making faithfulness relatively straightforward to evaluate. Additionally, summarization is commonly employed in fields such as news, medicine, and law, where reliability holds particular importance. Besides summarization, other tasks, such as text simplification, NMT, and dialogue generation, are also included in this section. Some evaluation methods concentrate on a single task, while others can be extended to similar ones.
3.1 With-reference Evaluators
For the summarization task, the approaches start with statistical methods. $\mathbf{Fact_{acc}}$ (Goodrich et al., 2019) uses information extraction models to measure factual overlap. The facts are pre-defined schema as triplet tuples (entity-relation-entity). Then the score is calculated by comparing the schema between ground-truths and generated summaries. Later, Kryscinski et al. (2020) propose a weakly-supervised model FactCC for verifying factual consistency of generated summaries. FactCC uses NLI learning structure to do fact classification. Additionally, the other version FactCCX can highlight spans in the source document that contain the support or opposition for the claim. Besides utilising entity-level facts, DAE (Goyal & Durrett, 2020) points out the dependency-level factual error, and designs a new entailment model. It views dependency arcs as semantic units and applies a 2-class classification task with automatically generated labels $\in\{entailed,non-entailed\}$, where entailing means the semantic relationship it implies between a dependency arc head and child is entailed by the ground-truth sentence, and vice versa.
Text simplification tasks aim to ”translate” texts to a more readable version without altering their meaning. Maskeval (Liu et al., 2022) proposed an automatic evaluator for summarization and simplification tasks. These two tasks share some nature in common, which is they both have long input and output text that need to be loyal to the source. It gets weights from masked LM for every token and multiplies the weight and the Exact Matching (EM) score to get the final score. The experiments show that weights trained on summarization can improve the metric’s performance on simplification, which again confirms the similarity between the two tasks.
Hallucinations in NMT have also received some attention so far (Lee et al., 2018; Raunak et al., 2021).  Guerreiro et al. (2023) proposes a hallucination detection method using glass-box uncertainty-based detectors. This paper states that hallucination in NMT is so rare that like a needle in a haystack.  Dale et al. (2023) also provide a detection method by computing the percentage of the source contribution to the generated translation. It includes reference-based oracles, and both internal and external methods to measure the hallucination percentage.
The above approaches provide a promising starting, as they explore existing datasets and deliver feasible solutions. For long-input tasks such as summarization and simplification, they can use label information, with manual correction making it more accurate. However, a notable drawback is the requirement for annotation, which is costly and labour-intensive, and may also need cross-checking. Additionally, for NMT, the source text serves as the reference. Distinguishing between normal translation errors and hallucinations presents significant challenges for models as well.
3.2 Reference-free Evaluators
Reference-free means no direct ground truth are provided while evaluating. For abstractive summarization, Question Generation and Question Answer (QA-QG) is a possible pipeline to deal with it. FEQA (Durmus et al., 2020) is an evaluation framework that first extracts question-answer pairs from the summary and then uses an off-the-shelf QA model to extract answers from the document to check if the answers are consistent. QAGS (Wang et al., 2020) follows the framework of FEQA. It uses a fine-tuned BART on NewsQA (Trischler et al., 2017) for QG, which uses cloze-style questions and uses a fine-tuned BERT on SQuAD2.0 (Rajpurkar et al., 2018) for QA. QuestEval (Scialom et al., 2021) is a QA-QG pipeline as well. What’s different is that it lays out a comprehensive evaluation of four dimensions (consistency, coherence, fluency, and relevance). It borrows the QAGS framework for computing the precision score for the summary from the source document and uses SummaQA (Scialom et al., 2019) to compute the recall score, where a question weighted applies. QAFactEval (Fabbri et al., 2022) analyzes all components of the QA-based metric pipeline before, including answer selection, question generation, question answering, answer overlap evaluation, and question filtering. Among them, QAFactEval picks the best components through sound experiments and has made a more powerful combination. To avoid unanswerable-question noise, it also applies a $IsAnsweredSummFilter$ to score and filter the questions. Furthermore, the experiments show that entailment and QA-based metrics offer complementary signals that the combination can achieve better performance. MQAG (Manakul et al., 2023a) measures the consistency between summary and source document by computing the statistical distance of answers over automatically generated multiple-choice questions. The multiple-choice system allows for a more abstract understanding of information and makes it convenient for standard information-theoretic measures. Counterfactual estimation can also be used to evaluate factual consistency. CoCo (Xie et al., 2021) calculates the score by measuring the causal effects from source documents to generated summaries. It can also be generalized to multi-level granularity and other tasks.
Extractive summarization just extracts sentences or sentence spans from source documents. ExtEval (Zhang et al., 2023b) is the first to reveal that extractive summarization also can not guarantee faithfulness. It defines five types of broad unfaithfulness problems which the entailment model can not cover all. Additionally, ExtEval uses four sub-metrics summation to get the unfaithful score.
For Dialog Generation, $\mathbf{Q^{2}}$ (Honovich et al., 2021) is a evaluation method for neural knowledge-grounded generative models. It incorporates QG, QA, and NLI models to score the model’s output to reflect its consistency with its grounding source knowledge. Specifically, it generates questions from named entities and gets answers from the knowledge. Finally, it measures the answer span similarity using an NLI model.
For multiple tasks, AlignScore (Zha et al., 2023) can give a unified function for 7 tasks consistency evaluation. It breaks the context into chunks and the claim into sentences to calculate the alignment score. WeCheck (Wu et al., 2023) finds previous metrics may fail to generalize to real generated samples due to the gap between training and testing and the scarcity of labelled data. It provides a weak supervision learning paradigm based on previous metrics and uses a mixture NLI datasets to warm up and do a noise-aware fine-tuning to get the target metric model.
In comparison to with-reference methods, these approaches have greater generality. It is noticeable that QG-QA models take a large portion. Because, without references, NLI models require accurate extraction of evidence from the source text, which is challenging. Essentially, the QG-QA framework is a flexible approach that extracts and compares information from generated output and source input. When comparing, some methods also incorporate entailment models for judgment. However, the QG and QA steps may introduce considerable noise, which can interfere with the results.
3.3 Datasets and Benchmarks
In addition to the evaluation models above, other studies have proposed helpful datasets and benchmarks to measure hallucination across tasks and to assess the efficacy of evaluators.
While building models, some works also release annotated datasets as well. For the summarization task,  Maynez et al. (2020) find severe extrinsic hallucination in the reference of XSum dataset, and label XSumFaith with hallucination type (intrinsic or extrinsic) and hallucination span position. CoGenSumm (Falke et al., 2019) is collected from the validation set of CNN/DM and is used to rerank generated summaries. From both CNN/DM and XSum, QAGS (Wang et al., 2020) annotates each sentence with a binary choice based on whether the sentence is faithful to the article. However, the binary factuality of texts can be difficult to determine. FRANK (Pagnoni et al., 2021) collect annotation based on a more fine-grained defined typology of factual errors. Polytope (Huang et al., 2020) provides both SF and WF annotations to measure both extractive and abstractive summarization.  Devaraj et al. (2022) label 3 types of factual errors for data collected from Newsela (Xu et al., 2015) and Wikilarge (Zhang & Lapata, 2017) for text simplification. In dialogue generation, Dialogue NLI (Welleck et al., 2018) provides three-type labels for NLI models of the entailment of sentence pairs.
Given the large amount of evaluation metrics proposed, some benchmarks are designed to reevaluate these metrics. SummEval (Fabbri et al., 2021), SummaC (Laban et al., 2022), Dialsummeval (Gao & Wan, 2022), and AGGREFACT (Tang et al., 2023) focus on summarization hallucination evaluation or detection. Wizard of Wikipedia (Dinan et al., 2018), CI-ToD (Qin et al., 2021), BEGIN (Dziri et al., 2022b), FaithDial (Dziri et al., 2022a) and TopicalChat (Gopalakrishnan et al., 2023) can help measure consistency evaluators on dialogue generation. Beyond a single task, TRUE (Honovich et al., 2022) and BEAMetrics (Scialom & Hill, 2021) can evaluate metrics across a series of text generation tasks.
4 Evaluation Methods After LLMs
With the rapid development of LLMs and their human-like performance, LLMs possess instruction-tuned LLMs, like ChatGPT (OpenAI, 2022) or GPT-4 (Achiam et al., 2023), which offer an advantage that can take instruction in the natural language as human experts when evaluating. For SF and WF in LLM scenarios, it is mainly presented in the form of context consistency or factual imagination. Some works use LLMs to evaluate specific tasks, while other works evaluate the hallucination level of LLMs themselves. In this section, we will elaborate on both aspects.
4.1 LLMs as Tool
Recent LLMs are powerful tools that can be prompted by instructions for evaluation. The criteria can be designed by task, so it is more flexible and can evaluate from multiple aspects.
Following the works of Sec. 3, SCALE (Lattimer et al., 2023) focuses on long-form dialogue and uses LLM as an NLI model to generate scores regarding to the prompts and generated texts. It breaks long source documents into chunks of arbitrary length to measure the supportive level of the source text span and also provides a long-form dialogue dataset ScreenEval from TV scripts and summaries pulled from the SummScreen.  Chen et al. (2023b) leverage prompting methods to evaluate summarization task. It designs chain-of-thoughts (CoT) and sentence-by-sentence prompts on few-shot and zero-shot scenarios and validates on a large range of summarization datasets.
Expanding to a larger scope of tasks, GPTScore (Fu et al., 2023) and G-Eval (Liu et al., 2023) both provide multi-faceted evaluation frameworks including consistency. GPTScore (Fu et al., 2023) covers a range of tasks including dialogue response generation, summarization, data-to-text, and NMT. When evaluating different tasks, normally, LLMs will be fed with the task introduction and the evaluation criteria as a prompt. GPTScore uses LLMs till GPT-3 (Brown et al., 2020) and can customize evaluation for different tasks. G-Eval (Liu et al., 2023) is backboned with GPT-4, using CoT and a form-filling paradigm. When evaluating, it is asked to generate a CoT of evaluation steps formatted as a form. Then the final score is the summation of weighted scores with predefined weight. The experiment results on summarization and dialogue generation showed that CoT can help GPT-4 better understand the evaluation criteria. Other than these two works, Wang et al. (2023b) studies the feasibility of ChatGPT as a good evaluator for three tasks. This paper shows us a way to build evaluators based on ChatGPT with or without reference. It explores both direct assessment and one-to-five stars ranking processes. ChainPoll (Friel & Sanyal, 2023) also uses ChatGPT to build a hallucination detection method with the CoT paradigm and considers it to be a binary classification problem.
Using LLMs for consistency evaluation generally results in better human correlation, which is an important direction for the development of evaluators. Human evaluation can be expensive, especially the experts who hold domain specialities. However, the performance of the above tools still is sensitive to the LLM’s capacity and the design of the prompts.
4.2 LLMs as Object
LLMs have a unique form, which themselves are dialogue systems and also can involve NLG tasks in the conversation. To implement evaluation, due to the flexible form and nonuniform test datasets, many works focus on proposing their own datasets, evaluation indicators, and establishing benchmarks. Among them, some studies assess LLM’s grasp of general facts and also updated world knowledge. Additionally, measuring hallucinations in long-context/generation and reasoning scenarios is also challenging, as is assessing LLM’s performance in specific domains and non-English contexts. Here we present recent evaluators and benchmarks for evaluating the hallucination level of LLMs from these aspects.
General Fact:
In a pipeline of fact extraction, evidence confirmation and fact consistency verification, FacTool and UFO both show competitive evaluation frameworks for hallucination. FacTool (Chern et al., 2023) is a tool-augmented framework that includes five main components: claim extraction, query generation, tool querying, evidence collection, and agreement verification. It includes multiple tasks and multiple domains, which makes it a relatively holistic evaluation method. UFO (Huang et al., 2024) provides a factuality verification method with multiple fact sources including human-written evidence, reference documents, search engine results, and LLM knowledge. The process has three modules: fact unit extraction, fact source verification, and fact consistency discrimination. It extracts entity-centric question-answer pairs as fact units and checks if the fact is consistent with the source. CONNER (Chen et al., 2023a) also relies on external evidence to test LLMs the world knowledge they possess. It performs knowledge-intensitive tasks, including open-domain question answering and knowledge-grounded dialogue, as intrinsic evaluation.
Utilizing LLM’s inherent capabilities for consistency checking is also a viable approach. SelfCheckGPT (Manakul et al., 2023b) is motivated by the simple idea that LLM tends to have self-consistency if it possesses certain knowledge. It uses the scheme of black-box zero-resource hallucination detection and compares multiple sampled responses then measuring consistency. SelfCheckGPT sentence-level score is computed by incorporating previous metrics, such as BertScore, question answer, and LLM prompt. Rather than self-consistency, ${}\mathbf{SAC^{3}}$ (Zhang et al., 2023a) performs cross-checking hallucination detection based on sampling. The cross-checking period includes question-level and model-level. By leveraging semantically equivalent question perturbation and cross-model response consistency checking, it can effectively identify the consistency and factuality of the answers.
Benefiting from the open source of recent LLMs, it is possible to access the parameters and dense distribution. EigenScore (Chen et al., 2024) exploits dense semantic information of LLMs to measure semantic consistency and diversity. Based on the observation of the existing extreme features in the LLM penultimate layer, it also designs a method (Test Time Feature Clipping) to prevent LLMs generate overconfident hallucinations.
In addition to the aforementioned evaluation frameworks, many benchmarks have also been proposed to facilitate a more comprehensive evaluation of LLMs. TruthfulQA (Lin et al., 2022) is a benchmark to measure whether a language model is truthful enough. It brings up three concerns of the hallucination in the QA system: accidental misuse, blocking positive applications, and malicious misuse. This paper suggests that, for improving truthfulness, only scaling up models is less promising than finetuning using training objectives, and larger models were more informative despite being less truthful. HaluEval (Li et al., 2023) finds that ChatGPT is likely to generate unverifiable hallucinations and hard to identify the hallucinations in the generated text. Experiments suggest that extra knowledge and reasoning steps for LLMs can help alleviate hallucinations. FACTOR (Muhlgay et al., 2023) is a benchmark that can measure LLM factuality concerning a controlled set of facts in a generation setting, including three sub-benchmarks: Wiki-FACTOR, News-FACTOR and Expert-FACTOR. It follows the error types from FRANK (Pagnoni et al., 2021) performs a multi-choice factual evaluation task.
Furthermore, for evaluating the metrics themselves, RealHall (Friel & Sanyal, 2023) is a benchmark for evaluation methods and contains both closed- and open-domain hallucinations, which correspond to SF and WF. FELM (Zhao et al., 2024) also provides a meta-evaluation for metrics. It broadens the conventional understanding of world knowledge factuality to diverse domain data including world knowledge, science, math, recommendation, and reasoning. It uses ChatGPT to generate the responses for prompts and do segmentation, while humans are asked to label factuality and error reasons, error types, and references for each sample.
Fresh Fact:
The world is constantly changing. How can we determine whether LLMs possess dynamic knowledge? The following benchmarks concentrate on constructing time-sensitive datasets to enable the evaluation of LLMs’ capacity to incorporate up-to-date information. FreshQA (Vu et al., 2023) is a QA benchmark that can question current world knowledge and also with false premises to the LLMs. The questions are divided into four types according to the freshness referring to the ever-changing world, including never-, slow-, fast-changing, and false-premise questions, which test the LLMs from different aspects. KoLA (Yu et al., 2023) use both Wikipedia and continuously collected emerging news and novels to evaluate the world knowledge of LLMs. It develops a self-contrast metric by contrasting two completions generated by the same model and gets the similarity score. RealTimeQA (Kasai et al., 2024) designs an annotation framework and submission workflow to support the updated evaluation system. It tests on both open- and closed-book QA systems and shows the importance of QA systems to be keep-dated. ERBench (Oh et al., 2024) verified facts through entity-relationship (ER) model. It can construct questions using templates from updated ER databases and evaluates LLMs with binary or multi-choice QA tasks.
Long Context/Generation:
Despite recent advancements in LLMs enabling them to better handle long texts, hallucination evaluation in long context or generation remains a challenge. FACTSCORE (Min et al., 2023) focuses on long-form text generated by LLMs. It breaks the generated text into a series of fine-grained atomic facts by InstructGPT and then each atomic fact is assigned a binary label. The data is about biographies of people sampled from English Wikipedia. Rather than considering whether the fact is globally truthful, FACTSCORE refers to a given source of knowledge. In this way, it can avoid information conflictions between different sources. BAMBOO (Dong et al., 2023) provides a multi-task long context benchmark for LLMs which also can detect hallucination. It uses an entailment model on two synthetic datasets named SenHallu and AbsHallu to check the hallucination level of LLMs. RV (Yang et al., 2023) performs zero-resource hallucination detection by reverse validation. It focuses on passage-level detection rather than sentence-level, which holds greater practical significance. This method first constructs a question for the entity of user concern from LLM’s output, then uses it to query the LLM and compare the answers. This paper also publishes a benchmark PHD for the evaluation of passage-level detection methods.
Factual Reasoning:
To evaluate the factual reasoning abilities of LLMs, Laban et al. (2023) build a benchmark SUMMEDITS. This work provides a three-step protocol for inconsistency detection benchmark creation and implements it in a 10-domain benchmark. It shows current LLMs are not yet proficient at complex factual reasoning
Domain Specific:
Hallucinations generated by LLMs in the medical field can have serious consequences in healthcare applications. Therefore MedHalt (Pal et al., 2023) designs a pipeline to implement three fact-related tests: False Confidence Test (FCT), None of the Above Test (Nota), and Fake Questions Test (FQT). Through these tests, MedHalt can effectively identify and filter hallucinations within the medical domain.
Chinese Language:
Numerous Chinese LLMs have also emerged along with the trend and hallucination is also a crucial problem. ChineseFactEval (Wang et al., 2023a) conducts the hallucination evaluation in seven scenarios, covering areas where generative AI may be applied in daily life, including general knowledge, scientific, medical, legal, financial, mathematics, and modern Chinese history scenarios. UHGEval (Liang et al., 2023) provides a Chinese-language dataset for LLM hallucination evaluation, generated by Chinese LLMs in news domain. The evaluation method uses LLMs for discriminative evaluation (binary classification), selective evaluation (multiple-choice question), and generative evaluation (use existing lexical and semantic metrics). HalluQA (Cheng et al., 2023) is a benchmark categorizing hallucinations into imitative falsehoods and factual errors. Inspired by TruthfulQA, it summarizes the question patterns and combines them with Chinese culture to create a new dataset. It also instructs GPT-4 to act as a judger to get the consistency rate. The result shows this benchmark maintains a challenging benchmark while 18 out of 24 LLMs rate below 50%.
It should be noted that there is a small portion of overlap between Sec. 4.2 and Sec. 4.1, meaning that some approaches utilize LLMs to evaluate LLMs (Chern et al., 2023; Zhao et al., 2024) as well. For instance, due to its outstanding performance, GPT-4 can do annotation instead of human sometimes (Cheng et al., 2023). In contrast to evaluators before LLMs that primarily focus on SF, methods outlined in this section not only can address SF evaluation but also pay more attention to the WF (general fact) aspect. This is because LLMs enable open-domain QA and can even leverage search engines for retrieval. Thus, assessing fresh facts is also an important aspect. Though the latest LLMs can support long-text processing, challenges still persist in analyzing long contexts or generating long outputs, leading to hallucinations caused by long-distance discourse. Factual errors can also accumulate in reasoning or multi-hop question answering. Furthermore, in other application domains such as the medical field and non-English language, some preliminary explorations have been conducted, which call for more work participation.
5 Conclusion and Future Direction
Evaluating hallucination in NLG is crucial as it determines the direction and trend for developing a better model. Historically, most metrics are task-based, as defining the criteria for task performance is relatively straightforward. As NLP research gradually shifts to LLMs, researchers have proposed new demands and challenges. We summarize several aspects of hallucination evaluation that urgently need to be addressed in the future.
Comprehensive evaluation Traditional evaluation methods based on a specific task are no longer sufficient to LLMs. Researchers have already proposed many comprehensive evaluation dimensions for the general capabilities, such as common sense and reasoning. A comprehensive and unified benchmark specialized for hallucination in the future will greatly promote model iteration to address hallucination issues.
Differentiation between hallucination and error Hallucinations and text errors can be confusing sometimes. According to the traditional definition of hallucination (smooth but incorrect), any response from a large model that differs from the ground truth can be considered as hallucination, which is obviously unreasonable and can mislead researchers. Some works in NMT have already made some progress (Dale et al., 2023), and it’s better for future evaluation methods to accurately identify real hallucinations.
Interpretabity Identifying fact granularity and analyzing hallucination reasons can provide significant assistance in solving hallucination problems. Some reasoning methods have the potential to analyze the underlying causes of hallucinations and have better evaluation. Other observing aspects lie in the internal state of model generation (Su et al., 2024), model structures, and even data distributions, which provide more analytical perspectives.
Long context/generation Current exploration of LLMs in long-context scenarios mainly focuses on long inputs, such as using a complete contract or multi-turn dialogue history as input. It is crucial to address hallucinations caused by the model’s difficulty in understanding long inputs. However, on the other hand, hallucinations of the type caused by contradictions between the beginning and end of long outputs are also worth exploring (Wei et al., 2024), such as detecting whether a model-generated novel contains inconsistencies in character behavior.
More application views LLMs have more challenging hallucinations in certain scenarios. In the multi-lingual scenario, low-resource languages other than English may be more prone to hallucinations, and cross-lingual conversation can face misunderstanding hallucinations due to different cultural backgrounds. Moreover, the latest research focuses on expanding LLMs to areas such as autonomous agents, multimodality, and real-world applications, which bring about new types of hallucinations, such as code hallucination (Qian et al., 2023), tool hallucination, visual hallucination, and so on. Detecting such hallucinations is very interesting and worthwhile direction to explore.
Appendix A Fact-checking
Fact-checking or fact-verification task is another line of work that has been paid much attention. The fact-checking framework can be divided into three components: claim detection, evidence retrieval, and claim verification (Guo et al., 2022), which is a relatively mature pipeline. Distinct from the evaluation methods discussed before in this paper, it typically involves assessing the factual accuracy of individual claims, mostly focusing on their WF. Wikipedia is a commonly used source for world knowledge (Thorne et al., 2018; Schuster et al., 2021; Kamoi et al., 2023; Gupta et al., 2022; Schuster et al., 2021), not only for fact-checking, but also for factuality evaluation. Especially when extracting evidence from a specific source, the WF turns into SF, which also demonstrates the dialectical unity of WF and SF. Benefiting from LLMs, fact-checking can process longer and more complex texts with more confidence and efficiency (Wang et al., 2023c). Due to the nature of the fact-checking task, it can be seen as a WF evaluator for text generation with a binary (true/false) checker.
Appendix B Human Evaluation
For hallucination evaluation, human perspectives can play a pivotal role, providing datasets and establishing benchmarks for the development of automatic models. To build a human annotation framework, there are three aspects requiring consideration: 1) How to design error categories and unify guidelines for annotators; 2) How to ensure the reliability of human annotation; And 3) how to digitally present annotated results. Human evaluation can be time-consuming and is particularly inefficient for large-scale evaluations, but still is the most trustworthy way of model evaluation.
Appendix C Supplement Table
We provide a table of evaluators here, as in Table 1. For the Based-model column, it means the models that evaluators use to perform evaluation or generate synthetic data. Metric column means the calculating method to get the final score. ✓and ✗ in SF and WF columns mean the aspects that the evaluators focus on.
Era Name New Dataset Data Source Fact Definition Task Based-model Method Metric SF WF Before LLM with Ref $Fact_{acc}$ WikiFact Wikipedia, Wikidata KB Ent-rel -ent Summ Transformer NER, Binary Relation Classifier P, R, F1 FactCC FactCC CNN/DM, XSumFaith Sent Summ BERT NLI (2-class) Likelihood DAE DAE PARANMT50M Dependency Summ ELECTRA NLI (2-class) Likelihood Maskeval / CNN/DM, WikiLarge, ASSET Word Summ, Simp T5 Word Weighting Weighted Match Score  Guerreiro et al. (2023) Haystack WMT2018 DE-EN Text Span NMT Transformer Uncertainty Measure Avg. Similarity  Dale et al. (2023) / Haystack Text Span NMT Transformer Source Contribution Percentage Before LLM Ref Free FEQA FEQA CNN/DM, XSum Sent Span Summ BART (QG), BERT (QA) QG-QA Avg. F1 QAGS QAGS CNN/DM, XSum Ent, Noun Phrase Summ BART (QG), BERT (QA) QG-QA Avg. Similarity QuestEval / CNN/DM, Xsum Ent, Noun Summ T5 (QG, QA) QG-QA P, R, F1 QAFactEval / SummaC NP Chunk Summ BART (QG), ELECTRA (QA) QG-QA, NLI LERC MQAG / QAGS, XSumFaith, Podcast Assessment, SummEval Sent Span Summ T5 (QG), Longformer (QA) Multi-Choice QA Choice Statistical Distance CoCo / QAGS, SummEval Token, Span, Sent, Doc Summarization BART Counterfactual Estimation Avg. Likelihood Diff ExtEval ExtEval CNN/DM Discourse, Coreference, Sentiment Summ SpanBERT, RoBERTa Direct Prediction, Statistic Summation of Sub-scores $Q^{2}$ $Q^{2}$ WOW Sent Span Diag T5 (QG); Albert- Xlarge (QA); RoBERTa (NLI) QG-QA, NLI Likelihood AlignScore / 22 datasets from 7 tasks Sent NLI, QA, paraphrasing, Fact Verification, IR, Semantic Similarity, Summ RoberTa 3-way classification Likelihood WeCheck / TRUE Response Summ, Diag, Para, Fact Check DeBERTaV3 Weakly Supervised NLI Likelihood After LLM tool SCALE ScreenEval LLM, Human Sentence Long Diag Flan-T5 NLI Likelihood  Chen et al. (2023b) / SummEval, XSumFaith, Goyal21, CLIFF Response Summarization Flan-T5, code-davinci-002, text-davinci-003, ChatGPT, GPT-4 Vanilla/COT/ Sent-by-Sent Prompt Balanced Acc GPTScore / 37 datasets from 4 tasks Various Summ, Diag, NMT, D2T GPT-2, OPT, FLAN, GPT-3 Direct Assessment Direct Score G-Eval / SummEval, Topical-Chat, QAGS Response Summ, Diag GPT-4 COT, Form-filling Weighted Scores Wang et al. (2023b) / 5 datasets from 3 tasks Response Summ, D2T, Story Gen ChatGPT Direct Assessment, Rating Direct score ChainPoll RealHall-closed, RealHall-open COVID-QA, DROP, Open Ass prompts, TriviaQA Response Hallu Detect gpt-3.5-turbo Direct Assessment (2-class) Acc After LLM object FacTool / RoSE, FactPrompts, HumanEval, GSM-Hard, Self-instruct Claim, Response Knowledge-based QA, Code Generation, Math Reasoning, Sci-literature Review ChatGPT Claim Extraction, Query Generation, Tool Querying, Evidence Collection, Agreement Verification P, R, F1 UFO / NQ, HotpotQA, TruthfulQA, CNN/DM, Multi-News, MS MARCO Ent Open-domain/ Web Retrieval-based/ Expert-validated/ Retrieval-Augmented QA, News Fact Generation ChatGPT (gpt-3.5-turbo-1106) Fact Unit Extraction, Fact Source Verification, Fact Consistency Discrimination Avg. Sub-scores CONNER / NQ, WoW Sentence Open-domain QA, Knowledge-grounded Dialogue NLI-RoBERTa -large, ColBERTv2 3-way NLI Acc SelfCheckGPT SelfCheckGPT WikiBio Response Hallu Detect GPT-3 Direct Assessment, QA, BERTScore, NLI, Ngram AUC-PR $SAC^{3}$ / HotpotQA, NQ-open Response QA Generation gpt-3.5-turbo, Falcon-7b-instruct, Guanaco-33b Cross-checking, QA Pair Consistency AUROC EigenScore / CoQA, SQuAD, TriviaQA Natural Questions Inner State Open-book QA Closed-book QA LLaMA, OPT Semantic Consistency/ Diversity in Dense Embedding Space AUROC, PCC TruthfulQA TruthfulQA LLM, Human Response Multi-Choice QA, Generation GPT-3-175B Answer Match Score HaluEval Task-specific, General Alpaca, Task datasets ChatGPT Response QA, Summ, Knowledge- grounded Diag, Generation ChatGPT Direct Assessment Acc FACTOR Wiki-/News-/ Expert- FACTOR Wikipedia, Refin- edWeb, ExpertQA Sent Span Generation / FRANK Error Classification likelihood FELM FELM TruthfulQA, Quora, MMLU, GSM8K, ChatGPT, Human Text Span, Claim World Knowledge, Sci and Tech, Math, Writing and Recom- mendation, Reasoning Vicuna, ChatGPT, GPT4 Direct Assessment F1, Balanced Acc FreshQA Never/Slow Fast-changing, false-premise Human Response Generation / Answer Match Acc KoLA KoLA Wikipedia, Updated News and Novels Response Knowledge Memorization /Understanding/Applying /Creating / Self-contrast Answer Match Similarity RealTimeQA RealTimeQA CNN, THE WEEK, USA Today Response Multi-Choice QA, Generation GPT-3, T5 Answer Match Acc, EM, F1 ERBench ERBench Database 5 datasets from Kaggle Ent-Rel Binary/ Multiple -choice QA / Direct Assessment, String Matching Ans/Rat/ Ans-Rat Acc, Hallu Rate FactScore / Biographies in Wikipedia Atomic Fact Generation InstructGPT, ChatGPT, PerplexityAI Binary Classification P BAMBOO SenHallu, AbsHallu 10 datasets from 5 tasks Response Multi-choice tasks, Select tasks ChatGPT Answer Match P, R, F1 RV PHD Human Annotator Ent Generation ChatGPT Construct Query, Access Databases, Entity-Answer Match P, R, F1 SummEdits SummEdits 9 datasets from Summ task Span Summ, Reasoning gpt-3.5-turbo Seed summary verify, Summary edits, Annotation Balanced Acc MedHalt MedHalt MedMCQA, Medqa USMILE, Medqa (Taiwan), Headqa, PubMed Response Reasoning Hallu Test, Memory Hallu Test ChatGPT Answer Match Pointwise Score, Acc ChineseFactEval ChineseFactEval / Response Generation / FacTool, Human annotator Direct Score UHGEval UHGEval Chinese News Websites Keywords Generative/ Discriminative/ Selective Evaluator GPT-4 Answer Match, Similarity Acc, Similarity Score HalluQA HalluQA Human Response Generation GLM-130B, ChatGPT, GPT-4 Direct Assessment Non-hallu Rate
Table 1: Evaluators and Benchmarks meta information.

### v2

@@
-Can We Catch the Elephant? The Evolvement of Hallucination Evaluation on Natural Language Generation: A Survey
+Can We Catch the Elephant? A Survey of the Evolvement of Hallucination Evaluation on Natural Language Generation
 Abstract
-Hallucination in Natural Language Generation (NLG) is like the elephant in the room, obvious but often overlooked until recent achievements significantly improved the fluency and grammatical accuracy of generated text.
-For Large Language Models (LLMs), hallucinations can happen in various downstream tasks and casual conversations, which need accurate assessment to enhance reliability and safety.
-However, current studies on hallucination evaluation vary greatly, and people still find it difficult to sort out and select the most appropriate evaluation methods.
-Moreover, as NLP research gradually shifts to the domain of LLMs, it brings new challenges to this direction.
-This paper provides a comprehensive survey on the evolvement of hallucination evaluation methods, aiming to address three key aspects: 1) Diverse definitions and granularity of facts; 2) The categories of automatic evaluators and their applicability; 3) Unresolved issues and future directions.
+Hallucination in Natural Language Generation (NLG) is like the elephant in the room, obvious but often overlooked until recent achievements significantly improved the fluency and grammaticality of generated text.
+As the capabilities of text generation models have improved, researchers have begun to pay more attention to the phenomenon of hallucination.
+Despite significant progress in this field in recent years, the evaluation system for hallucination is complex and diverse, lacking clear organization.
+We are the first to comprehensively survey how various evaluation methods have evolved with the development of text generation models from three dimensions, including hallucinated fact granularity, evaluator design principles, and assessment facets.
+This survey aims to help researchers identify current limitations in hallucination evaluation and highlight future research directions.
+[…]
+[…]
+Can We Catch the Elephant? A Survey of the Evolvement of Hallucination Evaluation on Natural Language Generation
+Siya Qi, Yulan He, Zheng Yuan King’s College London, London, UK {siya.qi, yulan.he, zheng.yuan}@kcl.ac.uk
 1 Introduction
-The hallucination problem in Natural Language Generation (NLG) typically refers to situations where the generated text deviates from the facts in source or external knowledge.
-Like the elephant in the room, this problem has existed since the beginning of the development of NLG.
-However, in the era of Large Language Models (LLMs), with the rapid improvement of fluency and grammatical accuracy in text generation, this issue has gradually gained more attention.
-The evaluation of hallucination motivates the development of LLMs to be more reliable and safe.
-In this paper, we make a comprehensive survey on the hallucination evaluation methods and elaborate them according to the development process, which gives a broad view of current trends and future directions.
+Hallucination in the context of Natural Language Generation (NLG) typically refers to situations where the generated text is inconsistent with, or unsupported by, the input source or external knowledge.
+Like an elephant in the room, this problem has existed since the beginning of the development of NLG.
+As text generation models continue to evolve, technologies like Large Language Models (LLMs) have achieved grammatical correctness and fluency nearly indistinguishable from human writing.
+Consequently, hallucination has gradually surfaced and attracted increased attention.
+The evaluation of hallucination matters since it motivates the development of LLMs to be more reliable and safe.
+In this paper, we make a comprehensive survey on the evolution of hallucination evaluation methods through multiple perspectives as in Figure 1, which gives a broad view of current trends and future directions.
 The definition of hallucination varies a lot from work to work over time.
-In the initial stages, hallucination refers to the occurrence of semantically incorrect but grammatically viable content (Lee et al., 2018).
-This setup is typically rooted in traditional tasks such as Text Summarization and Neural Machine Translation (NMT), which often involve definite source information.
+In the initial stages, hallucination referred to the occurrence of semantically incorrect but grammatically viable content (Lee et al., 2018).
+This setup was typically rooted in traditional tasks such as Summarization and Neural Machine Translation (NMT), which often involve definite source information.
 The breakthrough came with the advent of LLMs like ChatGPT (OpenAI, 2022), crossing the barriers of conversational systems.
@@
 Many previous tasks can also be accomplished within a conversation by prompting LLMs with correct instructions (Ouyang et al., 2022).
-In the response of LLMs, hallucination occurs now and then, with a mix of faithful and factual problems.
-Faithfulness and factuality are two concepts that are closely related when describing hallucinations in the generated text and can be prone to be confusing in some circumstances.
-In this paper, we add prefixes to both of them for better understanding by introducing Source Faithfulness (SF) and World Factuality (WF)  [^1] In this paper, we also write SF as faithfulness/consistency, and WF as factuality..
+In the responses of LLMs, hallucination occurs now and then, with a mix of faithful and factual problems.
+Figure 1: The distribution of evaluators on various aspects, including the timeline, hallucinated fact granularity, method design principles, and assessment facets.
+Faithfulness and factuality are two concepts that are closely related when describing hallucinations in generated text and can be prone to confusion in some circumstances.
+In this paper, we add prefixes to both of them for better understanding by introducing Source Faithfulness (SF) and World Factuality (WF).
+[^1] In this paper, we also write SF as faithfulness/consistency, and WF as factuality.
 SF measures the degree to which the generated output accurately reflects and is consistent with the source input.
 It ensures that the output stays true to the information provided in the source.
-SF has a more limited scope, as there are specific source documents that can be used to substantiate and verify the generated text.
+SF has a limited scope, as there are specific sources that can be used to substantiate and verify the generated text.
 WF, on the other hand, assesses whether the generated output aligns with general world knowledge and facts.
-WF is a more expansive and challenging problem as it goes beyond the specific source and considers the broader context of common sense and established knowledge, which are more difficult to comprehensively collect and encode.
-Recent studies have recognised the critical importance of addressing and improving the SF and WF of generated text.
-Whether in existing NLG tasks or the more practical scenarios in the LLMs era, SF and WF problems manifest in different forms.
-For example, in the NMT task, generated translations detached from the source text are considered factually incorrect (Dale et al., 2023).
-In the summarization task, generated content usually should belong to the source document to be faithful, but some also argue that certain hallucinations can align with external facts (Dong et al., 2022).
-In the dialogue task, it is important to balance the hallucination and user experience (Santhanam et al., 2021).
-In LLM downstream tasks, the issue is often related to the prompt, which is designed for different scenarios and to retrieve the knowledge that existed in LLM.
-Apart from these, LLM faces unique difficulties, such as updated world information and false-premise questions (Vu et al., 2023; Kasai et al., 2024).
-Hallucinations are prevalent across various tasks, as it is not a problem that has been completely defined, yet it intuitively pertains to binary correctness.
-While humans can readily perceive some hallucinations, the model often struggles to do so.
+WF is a more expansive and challenging problem as it goes beyond the specific source and considers the broader context of common sense and established knowledge, which is more difficult to collect and encode comprehensively.
+Recent studies have recognized the critical importance of addressing and improving the SF and WF of generated text.
+Assessing from SF or WF aspects means the evaluators refer to different source information, which is closely tied to specific tasks.
+For example, in NMT, generated translations detached from the source text are considered factually incorrect (Dale et al., 2023).
+In summarization, summaries usually should be faithful to the source document, but some also argue that certain hallucinations can align with external facts (Dong et al., 2022).
+In dialogue generation, it is important to balance hallucination and user experience (Santhanam et al., 2021).
+In LLM downstream tasks, the issue is often related to the prompts, which are designed for different scenarios and to retrieve the knowledge that exists in LLMs.
+Apart from these, LLMs face unique difficulties, such as updating world information and handling false-premise questions (Vu et al., 2023; Kasai et al., 2024).
+Hallucinations are prevalent across various tasks, as the problem has not been completely defined, yet it intuitively pertains to binary correctness.
+While humans can readily perceive some hallucinations, models often struggle to do so.
 Moreover, in the current development of LLMs regarding reasoning, planning, task adaptability, etc., hallucination evaluation is an issue that requires careful consideration.
-Previous works have some introductions on benchmarks for LLM hallucinations (Huang et al., 2023; Zhang et al., 2023c; Ji et al., 2023; Huang et al., 2021), but they do not include evaluators before LLMs and more research after, thus resulting in a lack of comparison between developments before and after.
-Also, they do not specifically focus on evaluators, therefore lacking in-depth analysis of details, as well as their respective strengths and weaknesses.
+Previous works have some introductions on benchmarks for LLM hallucinations (Huang et al., 2023; Zhang et al., 2023c; Ji et al., 2023; Huang et al., 2021), but they do not include evaluators before LLMs and more recent research, thus resulting in a lack of comparison between developments before and after.
+Also, they do not specifically focus on evaluators, therefore lacking in-depth analysis of details, as well as their strengths and weaknesses.
 Our contributions lie in three folds.
-Firstly, this paper organizes automatic evaluation methods for hallucination in the evolvement timeline, including traditional approaches as well as novel methods employing or assessing LLMs, and conducts analytical comparisons.
-Secondly, an applicable classification method is proposed, categorizing evaluation methods from a practical perspective based on the scenario.
-Lastly, recommendations for future hallucination evaluations are provided from multiple perspectives.
-We also give a brief introduction of related fact-checking and human evaluation works in Appendix A and B.
-2 Definition and Taxonomy
-To understand how to evaluate hallucination in text generation, we first need to define fact.
-In previous works, fact granularity is a way to define facts and serves as the foundational basis for assessing hallucination.
-Fact error classification also can help provide a solid and comprehensive perspective for evaluation.
+Firstly, this paper organizes automatic evaluation methods for hallucination [^2] We also give a brief introduction of related fact-checking and human evaluation works in Appendix A and B.
+in the evolution timeline, including traditional approaches as well as novel methods employing or assessing LLMs, and conducts analytical comparisons.
+Secondly, this paper reviews various evaluation methods from three perspectives: hallucinated fact granularity, evaluator design principles, and assessment facets, highlighting the intrinsic connections and development trajectories among these methods.
+Finally, we summarize the current state of research on hallucination evaluation, outlining existing issues and potential future directions.
+2 Fact Definition
+To evaluate hallucination in text generation, we first need to define what constitutes a fact.
+Previous works have evaluated hallucination by considering different levels of fact granularities and categorizing fact errors into various types.
 2.1 Fact Granularity
-The definition of fact granularity varies, lacking an agreed-upon labelling scheme (Honovich et al., 2022).
-Moreover, in generated text, many contents are difficult to have clear definitions.
-Among previous works, fact granularity can be categorized into the following types.
-Token/Word At the token/word level, entities and nouns are the most frequently used granularity, primarily due to their intuitively informative and relative ease of acquisition and comparison.
-For example, some summarization models tend to forget entities (Maynez et al., 2020; Nan et al., 2021), making up entities that never appeared in the source document.
-Span Spans require more annotation than individual tokens or words.
-Mostly, they are entity-centric, relation-centric, or event-centric, which are closely related to dependency parsing (Goyal & Durrett, 2020; 2021).
+The definition of fact granularity varies, lacking an agreed-upon labeling scheme (Honovich et al., 2022), and many elements in generated text are difficult to define clearly.
+Previous works categorize fact granularity into the following types.
+Figure 2: Source Faithful Error and World Factual Error examples.
+Token/Word: At the token/word level, entities and nouns are the most frequently used granularities, primarily due to their intuitively informative nature and relative ease of acquisition and comparison.
+For example, some summarization models tend to forget (Maynez et al., 2020; Nan et al., 2021) or make up entities that never appeared in the source document.
+Span: Spans require more annotation than individual tokens or words.
+Mostly, they are entity-centric, relation-centric, or event-centric, closely related to dependency parsing (Goyal and Durrett, 2020, 2021).
 In evaluators, spans can be used to generate questions (Durmus et al., 2020; Honovich et al., 2021) or check entailment (Kamoi et al., 2023).
-Sentence A higher level is sentence, which is coarsely classified whether it contains hallucination or not (Wang et al., 2020; Pagnoni et al., 2021).
-Sentence-level facts can help obtain an overall perception of the factual content within the sentence.
-Passage Passage-level facts can contribute to an overall assessment of the generated text (Fabbri et al., 2021).
-This encompasses discourse fact connections (Zhang et al., 2023b) and an estimation of passage themes.
-Moreover, this is crucial for future assessments of the factual accuracy in generating longer texts.
+Sentence: A higher level of granularity is the sentence, which is coarsely classified as containing hallucination or not (Wang et al., 2020; Pagnoni et al., 2021).
+Sentence-level facts provide an overall perception of the factual content within the sentence, allowing for a more holistic assessment.
+Passage: Passage-level facts contribute to an overall assessment of the generated text (Fabbri et al., 2021), encompassing discourse fact connections (Zhang et al., 2023b) and an estimation of passage themes.
+This level of granularity is crucial for evaluating longer texts.
+Figure 1 illustrates that evaluation methods before the advent of LLMs mostly relied on text segments, whereas many methods developed after LLMs frequently conduct direct assessments of the entire input text.
 2.2 Fact Error Type
-The taxonomies of fact errors vary subjecting to scenarios, some categorize them by the information source as intrinsic and extrinsic hallucinations (Maynez et al., 2020; Goyal & Durrett, 2021; Cao & Wang, 2021), For faithful error types,  Goyal & Durrett (2021) designs entity-, event-, and noun phrase-related errors.
-Frank (Pagnoni et al., 2021) categorizes with a more fine-grained scale, including PredE, EntE, CircE, CorefE, LinkE, OutE and GramE.
-We summarize that such errors can be categorized into Source Faithful Error and World Factual Error, in which way it can cover more NLG tasks including traditional ones and LLM scenarios.
-SF and WF errors are independent as shown in Figure 1.
-Figure 1: Source Faithful Error and World Factual Error examples.
+The taxonomies of fact errors vary depending on the scenarios.
+Some categorize them by the information source as intrinsic and extrinsic hallucinations (Maynez et al., 2020; Goyal and Durrett, 2021; Cao and Wang, 2021).
+For faithful error types, Goyal and Durrett (2021) designs entity-, event-, and noun phrase-related errors, while Pagnoni et al.
+(2021) categorize errors on a more fine-grained scale, including PredE, EntE, CircE, CorefE, LinkE, OutE and GramE.
+We propose categorizing errors into SF Error and WF Error, which can cover a wider range of NLG task, including both traditional ones and LLM scenarios.
+SF and WF errors can be assessed independently, as illustrated in Figure 2.
 3 Evaluation Methods Before LLMs
-This section contains the automatic evaluation methods for hallucinations in NLG before LLM, which differ from other evaluators that focus on text similarity or grammar correctness.
-Specifically, the focus mostly lies on the SF aspect of text generation.
-While previous works rely on the reference to train the models, the reference itself can suffer a lot of factual errors (Maynez et al., 2020), which hinder the model training from learning faithfulness.
-Also due to the costly but error-prone reference, many works study reference-free evaluation methods, which only require the input of source documents.
-To introduce the approaches, we will elaborate from evaluators to benchmarks.
-It is noteworthy that a significant portion of these evaluation methods is dedicated to assessing summarization tasks, due to which has a fixed source document as an information reference, making faithfulness relatively straightforward to evaluate.
-Additionally, summarization is commonly employed in fields such as news, medicine, and law, where reliability holds particular importance.
-Besides summarization, other tasks, such as text simplification, NMT, and dialogue generation, are also included in this section.
+This section discusses automatic evaluation methods for hallucinations in NLG before the era of LLMs.
+These methods differ from those that focus on text similarity (Lin and Hovy, 2003) or grammatical correctness (Bryant et al., 2017), as their primary focus lies on the SF aspect of text generation.
+Previous works rely on references to train models, but these references can contain factual errors (Maynez et al., 2020), hindering the model’s ability to learn faithfulness.
+Due to the costly and error-prone nature of references, many works study reference-free evaluation methods that only require the input of source documents.
+A significant portion of these evaluation methods is dedicated to assessing the summarization task, which has a fixed source document as an information reference, making faithfulness relatively straightforward to evaluate.
+Summarization is commonly employed in fields such as news, medicine, and law, where reliability is of particular importance.
+In addition to summarization, this section covers tasks such as text simplification, NMT, and dialogue generation.
 Some evaluation methods concentrate on a single task, while others can be extended to similar ones.
-3.1 With-reference Evaluators
-For the summarization task, the approaches start with statistical methods.
-$\mathbf{Fact_{acc}}$ (Goodrich et al., 2019) uses information extraction models to measure factual overlap.
-The facts are pre-defined schema as triplet tuples (entity-relation-entity).
-Then the score is calculated by comparing the schema between ground-truths and generated summaries.
-Later, Kryscinski et al.
-(2020) propose a weakly-supervised model FactCC for verifying factual consistency of generated summaries.
-FactCC uses NLI learning structure to do fact classification.
-Additionally, the other version FactCCX can highlight spans in the source document that contain the support or opposition for the claim.
-Besides utilising entity-level facts, DAE (Goyal & Durrett, 2020) points out the dependency-level factual error, and designs a new entailment model.
-It views dependency arcs as semantic units and applies a 2-class classification task with automatically generated labels $\in\{entailed,non-entailed\}$, where entailing means the semantic relationship it implies between a dependency arc head and child is entailed by the ground-truth sentence, and vice versa.
-Text simplification tasks aim to ”translate” texts to a more readable version without altering their meaning.
-Maskeval (Liu et al., 2022) proposed an automatic evaluator for summarization and simplification tasks.
-These two tasks share some nature in common, which is they both have long input and output text that need to be loyal to the source.
-It gets weights from masked LM for every token and multiplies the weight and the Exact Matching (EM) score to get the final score.
-The experiments show that weights trained on summarization can improve the metric’s performance on simplification, which again confirms the similarity between the two tasks.
-Hallucinations in NMT have also received some attention so far (Lee et al., 2018; Raunak et al., 2021).
+In the following subsections, we will introduce the approaches by elaborating on evaluators and benchmarks.
+3.1 Reference-based Evaluators
+Reference-based evaluators compare generated text with ground-truth references to assess factual consistency.
+For summarization, early approaches start with statistical methods.
+$\mathbf{Fact_{acc}}$ (Goodrich et al., 2019) uses information extraction models to measure factual overlap based on pre-defined schemas.
+FactCC and FactCCX Kryscinski et al.
+(2020) use Natural Language Inference (NLI) models for fact classification and span highlighting.
+Besides utilising entity-level facts, DAE (Goyal and Durrett, 2020) focuses on dependency-level factual error using entailment models.
+Text simplification aims to “translate” texts to a more readable version without altering their meaning.
+Maskeval (Liu et al., 2022) evaluates both summarization and simplification by combining masked LM weights with Exact Matching (EM) scores, exploiting the similarity between the two tasks.
+Hallucinations in NMT have received attention as well (Lee et al., 2018; Raunak et al., 2021).
+Hallucination in NMT is rare, like finding a needle in a haystack.
+It is important to identify the hallucinations with the errors.
 Guerreiro et al.
-(2023) proposes a hallucination detection method using glass-box uncertainty-based detectors.
-This paper states that hallucination in NMT is so rare that like a needle in a haystack.
-Dale et al.
-(2023) also provide a detection method by computing the percentage of the source contribution to the generated translation.
-It includes reference-based oracles, and both internal and external methods to measure the hallucination percentage.
-The above approaches provide a promising starting, as they explore existing datasets and deliver feasible solutions.
+(2023) propose a glass-box uncertainty-based detection method, while  Dale et al.
+(2023) measure hallucination by computing the percentage of source contribution to the generated translation.
+These approaches provide promising solutions, as they explore existing datasets and deliver feasible solutions.
 For long-input tasks such as summarization and simplification, they can use label information, with manual correction making it more accurate.
-However, a notable drawback is the requirement for annotation, which is costly and labour-intensive, and may also need cross-checking.
-Additionally, for NMT, the source text serves as the reference.
-Distinguishing between normal translation errors and hallucinations presents significant challenges for models as well.
+However, a notable drawback is the requirement for annotation, which is costly and labor-intensive, and may also need cross-checking.
+Additionally, distinguishing between normal translation errors and hallucinations in NMT presents significant challenges.
 3.2 Reference-free Evaluators
-Reference-free means no direct ground truth are provided while evaluating.
-For abstractive summarization, Question Generation and Question Answer (QA-QG) is a possible pipeline to deal with it.
-FEQA (Durmus et al., 2020) is an evaluation framework that first extracts question-answer pairs from the summary and then uses an off-the-shelf QA model to extract answers from the document to check if the answers are consistent.
-QAGS (Wang et al., 2020) follows the framework of FEQA.
-It uses a fine-tuned BART on NewsQA (Trischler et al., 2017) for QG, which uses cloze-style questions and uses a fine-tuned BERT on SQuAD2.0 (Rajpurkar et al., 2018) for QA.
-QuestEval (Scialom et al., 2021) is a QA-QG pipeline as well.
-What’s different is that it lays out a comprehensive evaluation of four dimensions (consistency, coherence, fluency, and relevance).
-It borrows the QAGS framework for computing the precision score for the summary from the source document and uses SummaQA (Scialom et al., 2019) to compute the recall score, where a question weighted applies.
-QAFactEval (Fabbri et al., 2022) analyzes all components of the QA-based metric pipeline before, including answer selection, question generation, question answering, answer overlap evaluation, and question filtering.
-Among them, QAFactEval picks the best components through sound experiments and has made a more powerful combination.
-To avoid unanswerable-question noise, it also applies a $IsAnsweredSummFilter$ to score and filter the questions.
-Furthermore, the experiments show that entailment and QA-based metrics offer complementary signals that the combination can achieve better performance.
-MQAG (Manakul et al., 2023a) measures the consistency between summary and source document by computing the statistical distance of answers over automatically generated multiple-choice questions.
-The multiple-choice system allows for a more abstract understanding of information and makes it convenient for standard information-theoretic measures.
-Counterfactual estimation can also be used to evaluate factual consistency.
-CoCo (Xie et al., 2021) calculates the score by measuring the causal effects from source documents to generated summaries.
+Reference-free evaluation methods assess the quality of generated text without relying on direct ground truth.
+For abstractive summarization, the Question Generation and Question Answer (QG-QA) pipeline is a common approach.
+FEQA (Durmus et al., 2020), QAGS (Wang et al., 2020), and QuestEval (Scialom et al., 2021) follow a similar framework: extracting question-answer pairs from the summary and then using off-the-shelf QA models to check the consistency of answers extracted from the document.
+QuestEval also evaluates four dimensions beyond faithfulness and uses a weighter to identify important questions.
+QAFactEval (Fabbri et al., 2022) analyzes and combines the best components of the QA-based metric pipeline for improved performance, including answer selection, question generation, question answering, answer overlap evaluation, and question filtering.
+MQAG (Manakul et al., 2023a) computes the statistical distance of answers over automatically generated multiple-choice questions, allowing for a more abstract understanding of information.
+Counterfactual estimation, such as CoCo (Xie et al., 2021), measures the causal effects from source documents to generated summaries.
 It can also be generalized to multi-level granularity and other tasks.
-Extractive summarization just extracts sentences or sentence spans from source documents.
-ExtEval (Zhang et al., 2023b) is the first to reveal that extractive summarization also can not guarantee faithfulness.
-It defines five types of broad unfaithfulness problems which the entailment model can not cover all.
-Additionally, ExtEval uses four sub-metrics summation to get the unfaithful score.
-For Dialog Generation, $\mathbf{Q^{2}}$ (Honovich et al., 2021) is a evaluation method for neural knowledge-grounded generative models.
-It incorporates QG, QA, and NLI models to score the model’s output to reflect its consistency with its grounding source knowledge.
-Specifically, it generates questions from named entities and gets answers from the knowledge.
-Finally, it measures the answer span similarity using an NLI model.
+Other than QA-QG pipeline, FactGraph (Ribeiro et al., 2022) compares the generated summary with the input document through semantic graphs, thereby enhancing discourse comprehensibility.
+FactKB (Feng et al., 2023) introduces external triplet knowledge to augment the training data, and assess the consistency of news summarization.
+ExtEval (Zhang et al., 2023b) reveals that extractive summarization does not guarantee faithfulness, identifying five broad categories of unfaithfulness issues that the entailment model cannot fully address.
+For Dialog Generation, $\mathbf{Q^{2}}$ (Honovich et al., 2021) is a evaluation method for knowledge-grounded generative models.
+It incorporates QG, QA, and NLI models to measure the model’s output to reflect its consistency with its grounding source knowledge, and measures the answer span similarity using an NLI model.
+FactPush (Steen et al., 2023) provides a simple data augmentation method by adding phrases to sentences, which can improve the NLI model performance.
 For multiple tasks, AlignScore (Zha et al., 2023) can give a unified function for 7 tasks consistency evaluation.
 It breaks the context into chunks and the claim into sentences to calculate the alignment score.
-WeCheck (Wu et al., 2023) finds previous metrics may fail to generalize to real generated samples due to the gap between training and testing and the scarcity of labelled data.
-It provides a weak supervision learning paradigm based on previous metrics and uses a mixture NLI datasets to warm up and do a noise-aware fine-tuning to get the target metric model.
-In comparison to with-reference methods, these approaches have greater generality.
+WeCheck (Wu et al., 2023) provides a weak supervision learning paradigm based on previous metrics and uses a mixture NLI datasets to warm up and do a noise-aware fine-tuning to get the target metric model.
+Compared to reference-based methods, these approaches have greater generality.
 It is noticeable that QG-QA models take a large portion.
-Because, without references, NLI models require accurate extraction of evidence from the source text, which is challenging.
+Because, without references, it is challenging for NLI models to get accurate extraction of evidence from the source text.
 Essentially, the QG-QA framework is a flexible approach that extracts and compares information from generated output and source input.
+However, the QG and QA steps may introduce considerable noise, which can interfere with the results.
 When comparing, some methods also incorporate entailment models for judgment.
-However, the QG and QA steps may introduce considerable noise, which can interfere with the results.
+It is showed that entailment and QA-based metrics offer complementary signals that the combination can achieve better performance (Fabbri et al., 2022).
 3.3 Datasets and Benchmarks
-In addition to the evaluation models above, other studies have proposed helpful datasets and benchmarks to measure hallucination across tasks and to assess the efficacy of evaluators.
-While building models, some works also release annotated datasets as well.
-For the summarization task,  Maynez et al.
+Alongside evaluation models, datasets and benchmarks have been proposed to measure hallucination across tasks and assess evaluator efficacy.
+For summarization,  Maynez et al.
 (2020) find severe extrinsic hallucination in the reference of XSum dataset, and label XSumFaith with hallucination type (intrinsic or extrinsic) and hallucination span position.
@@
 Polytope (Huang et al., 2020) provides both SF and WF annotations to measure both extractive and abstractive summarization.
+Falsesum (Utama et al., 2022) mimics intrinsic errors and extrinsic errors to augment summary datasets for NLI models.
 Devaraj et al.
-(2022) label 3 types of factual errors for data collected from Newsela (Xu et al., 2015) and Wikilarge (Zhang & Lapata, 2017) for text simplification.
+(2022) label 3 types of factual errors for data collected from Newsela (Xu et al., 2015) and Wikilarge (Zhang and Lapata, 2017) for text simplification.
 In dialogue generation, Dialogue NLI (Welleck et al., 2018) provides three-type labels for NLI models of the entailment of sentence pairs.
 Given the large amount of evaluation metrics proposed, some benchmarks are designed to reevaluate these metrics.
-SummEval (Fabbri et al., 2021), SummaC (Laban et al., 2022), Dialsummeval (Gao & Wan, 2022), and AGGREFACT (Tang et al., 2023) focus on summarization hallucination evaluation or detection.
+SummEval (Fabbri et al., 2021), SummaC (Laban et al., 2022), Dialsummeval (Gao and Wan, 2022), and AGGREFACT (Tang et al., 2023) focus on summarization hallucination evaluation or detection.
 Wizard of Wikipedia (Dinan et al., 2018), CI-ToD (Qin et al., 2021), BEGIN (Dziri et al., 2022b), FaithDial (Dziri et al., 2022a) and TopicalChat (Gopalakrishnan et al., 2023) can help measure consistency evaluators on dialogue generation.
-Beyond a single task, TRUE (Honovich et al., 2022) and BEAMetrics (Scialom & Hill, 2021) can evaluate metrics across a series of text generation tasks.
+Beyond a single task, TRUE (Honovich et al., 2022) and BEAMetrics (Scialom and Hill, 2021) can evaluate metrics across a series of text generation tasks.
 4 Evaluation Methods After LLMs
-With the rapid development of LLMs and their human-like performance, LLMs possess instruction-tuned LLMs, like ChatGPT (OpenAI, 2022) or GPT-4 (Achiam et al., 2023), which offer an advantage that can take instruction in the natural language as human experts when evaluating.
-For SF and WF in LLM scenarios, it is mainly presented in the form of context consistency or factual imagination.
-Some works use LLMs to evaluate specific tasks, while other works evaluate the hallucination level of LLMs themselves.
+With the rapid development of LLMs and their human-like performance, LLMs possess instruction-tuned LLMs, like ChatGPT (OpenAI, 2022) or GPT-4 (Achiam et al., 2023), which offer an advantage that can take instruction in natural language as human experts when evaluating.
+For SF and WF in LLM scenarios, they are mainly presented in the form of context consistency or factual imagination.
+Some works prompt LLMs as direct evaluators, while other works design a pipeline with multiple steps to do the assessment.
 In this section, we will elaborate on both aspects.
-4.1 LLMs as Tool
-Recent LLMs are powerful tools that can be prompted by instructions for evaluation. The criteria can be designed by task, so it is more flexible and can evaluate from multiple aspects.
+4.1 Direct Evaluation
+Recent LLMs are powerful tools that can be prompted by instructions for text evaluation.
+The criteria can be designed according to different tasks, so it is more flexible and can assess from multiple aspects.
+Meanwhile, to evaluate the level of LLMs’ hallucination, a variety of methods have been developed as well.
+This section explores these methods using direct assessment.
+These approaches include prompting the LLMs to generate responses or using a classifier to decide whether the input contains hallucinations.
+4.1.1 LLMs as tool
+In this section, we introduce the methods that leverage LLMs as tools to conduct hallucination evaluation.
 Following the works of Sec.
-3, SCALE (Lattimer et al., 2023) focuses on long-form dialogue and uses LLM as an NLI model to generate scores regarding to the prompts and generated texts.
-It breaks long source documents into chunks of arbitrary length to measure the supportive level of the source text span and also provides a long-form dialogue dataset ScreenEval from TV scripts and summaries pulled from the SummScreen.
+3, SCALE (Lattimer et al., 2023) focuses on long-form dialogue and uses the LLM as an NLI model to generate scores.
+It breaks long source documents into chunks to measure the supportive level of the source text span.
 Chen et al.
 (2023b) leverage prompting methods to evaluate summarization task.
-It designs chain-of-thoughts (CoT) and sentence-by-sentence prompts on few-shot and zero-shot scenarios and validates on a large range of summarization datasets.
+It designs chain-of-thoughts (CoT) and sentence-by-sentence prompts on few-shot and zero-shot scenarios.
+TrueTeacher (Gekhman et al., 2023) elaborates a data augmentation method, which uses LLM to label the generated summaries from CNN/DM, as well as multilingual data.
 Expanding to a larger scope of tasks, GPTScore (Fu et al., 2023) and G-Eval (Liu et al., 2023) both provide multi-faceted evaluation frameworks including consistency.
-GPTScore (Fu et al., 2023) covers a range of tasks including dialogue response generation, summarization, data-to-text, and NMT.
+GPTScore (Fu et al., 2023) covers a range of tasks including dialogue, summarization, data-to-text, and NMT.
 When evaluating different tasks, normally, LLMs will be fed with the task introduction and the evaluation criteria as a prompt.
-GPTScore uses LLMs till GPT-3 (Brown et al., 2020) and can customize evaluation for different tasks.
-G-Eval (Liu et al., 2023) is backboned with GPT-4, using CoT and a form-filling paradigm.
+G-Eval (Liu et al., 2023) uses CoT and a form-filling paradigm.
 When evaluating, it is asked to generate a CoT of evaluation steps formatted as a form.
-Then the final score is the summation of weighted scores with predefined weight.
-The experiment results on summarization and dialogue generation showed that CoT can help GPT-4 better understand the evaluation criteria.
+It shows that CoT can help GPT-4 better understand the evaluation criteria.
 Other than these two works, Wang et al.
-(2023b) studies the feasibility of ChatGPT as a good evaluator for three tasks.
+(2023b) study the feasibility of ChatGPT as a good evaluator for three tasks.
 This paper shows us a way to build evaluators based on ChatGPT with or without reference.
-It explores both direct assessment and one-to-five stars ranking processes.
-ChainPoll (Friel & Sanyal, 2023) also uses ChatGPT to build a hallucination detection method with the CoT paradigm and considers it to be a binary classification problem.
-Using LLMs for consistency evaluation generally results in better human correlation, which is an important direction for the development of evaluators.
-Human evaluation can be expensive, especially the experts who hold domain specialities.
-However, the performance of the above tools still is sensitive to the LLM’s capacity and the design of the prompts.
-4.2 LLMs as Object
-LLMs have a unique form, which themselves are dialogue systems and also can involve NLG tasks in the conversation.
-To implement evaluation, due to the flexible form and nonuniform test datasets, many works focus on proposing their own datasets, evaluation indicators, and establishing benchmarks.
-Among them, some studies assess LLM’s grasp of general facts and also updated world knowledge.
-Additionally, measuring hallucinations in long-context/generation and reasoning scenarios is also challenging, as is assessing LLM’s performance in specific domains and non-English contexts.
-Here we present recent evaluators and benchmarks for evaluating the hallucination level of LLMs from these aspects.
-General Fact:
-In a pipeline of fact extraction, evidence confirmation and fact consistency verification, FacTool and UFO both show competitive evaluation frameworks for hallucination.
-FacTool (Chern et al., 2023) is a tool-augmented framework that includes five main components: claim extraction, query generation, tool querying, evidence collection, and agreement verification.
-It includes multiple tasks and multiple domains, which makes it a relatively holistic evaluation method.
-UFO (Huang et al., 2024) provides a factuality verification method with multiple fact sources including human-written evidence, reference documents, search engine results, and LLM knowledge.
-The process has three modules: fact unit extraction, fact source verification, and fact consistency discrimination.
-It extracts entity-centric question-answer pairs as fact units and checks if the fact is consistent with the source.
-CONNER (Chen et al., 2023a) also relies on external evidence to test LLMs the world knowledge they possess.
-It performs knowledge-intensitive tasks, including open-domain question answering and knowledge-grounded dialogue, as intrinsic evaluation.
-Utilizing LLM’s inherent capabilities for consistency checking is also a viable approach.
-SelfCheckGPT (Manakul et al., 2023b) is motivated by the simple idea that LLM tends to have self-consistency if it possesses certain knowledge.
-It uses the scheme of black-box zero-resource hallucination detection and compares multiple sampled responses then measuring consistency.
-SelfCheckGPT sentence-level score is computed by incorporating previous metrics, such as BertScore, question answer, and LLM prompt.
-Rather than self-consistency, ${}\mathbf{SAC^{3}}$ (Zhang et al., 2023a) performs cross-checking hallucination detection based on sampling.
-The cross-checking period includes question-level and model-level.
-By leveraging semantically equivalent question perturbation and cross-model response consistency checking, it can effectively identify the consistency and factuality of the answers.
+ChainPoll (Friel and Sanyal, 2023) also uses ChatGPT to build a hallucination detection method with the CoT paradigm and considers it to be a binary classification problem.
+Using LLMs for evaluation is simple and convenient.
+Despite the increasing trust in LLMs as they grow in size and capability, determining how to apply LLMs effectively to ensure their stability and reliability remains an open question.
+4.1.2 LLMs as Object
+In this section, we present recent evaluators and benchmarks for evaluating the hallucination level of LLMs from many significant aspects.
+General Fact: TruthfulQA (Lin et al., 2022) is a benchmark to measure whether a language model is truthful enough.
+It brings up three concerns of the hallucination in the QA system: accidental misuse, blocking positive applications, and malicious misuse, and hedging is an acceptable solution for LLMs to be truthful in this work.
+HaluEval (Li et al., 2023) finds that ChatGPT is likely to generate unverifiable hallucinations and hard to identify the hallucinations in the generated text.
+Experiments suggest that extra knowledge and reasoning steps for LLMs can help alleviate hallucinations.
+FACTOR (Muhlgay et al., 2023) can measure LLM factuality concerning a controlled set of facts in a generation setting.
+It follows the error types from FRANK (Pagnoni et al., 2021) and performs a multi-choice factual evaluation task.
 Benefiting from the open source of recent LLMs, it is possible to access the parameters and dense distribution.
 EigenScore (Chen et al., 2024) exploits dense semantic information of LLMs to measure semantic consistency and diversity.
-Based on the observation of the existing extreme features in the LLM penultimate layer, it also designs a method (Test Time Feature Clipping) to prevent LLMs generate overconfident hallucinations.
-In addition to the aforementioned evaluation frameworks, many benchmarks have also been proposed to facilitate a more comprehensive evaluation of LLMs.
-TruthfulQA (Lin et al., 2022) is a benchmark to measure whether a language model is truthful enough.
-It brings up three concerns of the hallucination in the QA system: accidental misuse, blocking positive applications, and malicious misuse.
-This paper suggests that, for improving truthfulness, only scaling up models is less promising than finetuning using training objectives, and larger models were more informative despite being less truthful.
-HaluEval (Li et al., 2023) finds that ChatGPT is likely to generate unverifiable hallucinations and hard to identify the hallucinations in the generated text.
-Experiments suggest that extra knowledge and reasoning steps for LLMs can help alleviate hallucinations.
-FACTOR (Muhlgay et al., 2023) is a benchmark that can measure LLM factuality concerning a controlled set of facts in a generation setting, including three sub-benchmarks: Wiki-FACTOR, News-FACTOR and Expert-FACTOR.
-It follows the error types from FRANK (Pagnoni et al., 2021) performs a multi-choice factual evaluation task.
-Furthermore, for evaluating the metrics themselves, RealHall (Friel & Sanyal, 2023) is a benchmark for evaluation methods and contains both closed- and open-domain hallucinations, which correspond to SF and WF.
+Based on the observation of the existing extreme features in the LLM penultimate layer, it also use Test Time Feature Clipping to prevent LLMs generate overconfident hallucinations.
+Furthermore, for evaluating the metrics themselves, RealHall (Friel and Sanyal, 2023) is a benchmark for evaluation methods and contains both closed- and open-domain hallucinations, which correspond to SF and WF.
 FELM (Zhao et al., 2024) also provides a meta-evaluation for metrics.
@@
 It uses ChatGPT to generate the responses for prompts and do segmentation, while humans are asked to label factuality and error reasons, error types, and references for each sample.
-Fresh Fact:
-The world is constantly changing.
+Fresh Fact: The world is constantly changing.
 How can we determine whether LLMs possess dynamic knowledge?
 The following benchmarks concentrate on constructing time-sensitive datasets to enable the evaluation of LLMs’ capacity to incorporate up-to-date information.
-FreshQA (Vu et al., 2023) is a QA benchmark that can question current world knowledge and also with false premises to the LLMs.
-The questions are divided into four types according to the freshness referring to the ever-changing world, including never-, slow-, fast-changing, and false-premise questions, which test the LLMs from different aspects.
+FreshQA (Vu et al., 2023) is a QA benchmark that includes questions about current events and also inputs with false premises to the LLMs.
+The questions are divided into four types according to the freshness referring to the ever-changing world, including never-, slow-, fast-changing, and false-premise questions.
+RealTimeQA (Kasai et al., 2024) designs an annotation framework and submission workflow to support the updated evaluation system.
+It tests on both open- and closed-book QA systems and shows the importance of QA systems to be keep-dated.
+ERBench (Oh et al., 2024) verifies facts through entity-relationship (ER) model.
+It can construct questions using templates from updated ER databases and evaluates LLMs with binary or multi-choice QA tasks.
+Long Context/Generation: Despite recent advancements in LLMs enabling them to better handle long texts, hallucination evaluation in long context or generation remains a challenge.
+FACTSCORE (Min et al., 2023) focuses on long-form text generated by LLMs.
+It breaks the generated text into a series of fine-grained atomic facts and then each atomic fact is assigned a binary label.
+The data is about biographies of people sampled from English Wikipedia.
+It refers to a given source of knowledge, which avoids information conflictions between different sources, but challenges the comprehensiveness and accuracy of the knowledge data.
+BAMBOO (Dong et al., 2023) provides a multi-task long context benchmark for LLMs which also can detect hallucination, which uses an entailment model on two synthetic datasets.
+Domain Specific: Hallucinations generated by LLMs in the medical field can have serious consequences in healthcare applications.
+Therefore MedHalt (Pal et al., 2023) designs a pipeline to implement three fact-related tests: False Confidence Test (FCT), None of the Above Test (Nota), and Fake Questions Test (FQT).
+Through these tests, MedHalt can effectively identify and filter hallucinations within the medical domain.
+Chinese Language: Numerous Chinese LLMs have also emerged along with the trend and hallucination is also a crucial problem.
+ChineseFactEval (Wang et al., 2023a) conducts the evaluation in seven scenarios, covering areas where generative AI may be applied in daily life, and specifically includes the modern Chinese history scenario.
+UHGEval (Liang et al., 2023) provides a Chinese-language hallucination dataset, generated by Chinese LLMs in news domain.
+The evaluation method uses LLMs for discriminative, selective and generative evaluation.
+HalluQA (Cheng et al., 2023) is a benchmark categorizing hallucinations into imitative falsehoods and factual errors.
+Inspired by TruthfulQA, it summarizes the question patterns and combines them with Chinese culture.
+It should be noted that there is a small portion of overlap between Sec.
+4.1.2 and Sec.
+4.1.1, meaning that some approaches utilize LLMs to evaluate LLMs (Zhao et al., 2024) as well.
+For instance, due to its outstanding performance, GPT-4 can sometimes do annotation instead of human (Cheng et al., 2023).
+In contrast to evaluators before LLMs that primarily focus on SF, methods outlined in this section not only can address SF evaluation but also pay more attention to the WF (general fact) aspect, according to Figure 1.
+This is because LLMs enable open-domain QA and can even leverage search engines for retrieval.
+Thus, assessing fresh facts is also an important aspect.
+Though the latest LLMs can support long-text processing, challenges still persist in analyzing long contexts or generating long outputs, leading to hallucinations caused by long-distance discourse.
+How to catch the discourse-level errors while maintaining the ever-changing context is still challenging.
+Furthermore, in other application domains such as the medical field and non-English language, the collection of domain-specific data and the design of methodologies require further investigation, particularly in unique contexts such as medical statistics, legal terminology, and historical culture.
+4.2 Pipeline Evaluation
+Apart from utilizing LLMs for direct assessment, some evaluation methods do not rely on specific knowledge datasets or the pre-existing knowledge of LLMs.
+Instead, they leverage relevant evidence acquired through LLMs or search engines, benefitting from their enhanced search capabilities.
+Additionally, some methods involve the model generating answers multiple times to cross-check consistency to get the assessment outcome.
+In a pipeline of fact extraction, evidence confirmation and fact consistency verification, FacTool (Chern et al., 2023) and UFO (Huang et al., 2024) both show competitive evaluation frameworks for hallucination.
+FacTool is a tool-augmented framework that includes includes multiple tasks and domains, which makes it relatively holistic.
+UFO provides a factuality verification method with multiple fact sources including human-written evidence, reference documents, search engine results, and LLM knowledge.
+CONNER (Chen et al., 2023a) also relies on external evidence to test LLMs the world knowledge they possess.
+It performs knowledge-intensitive tasks, including open-domain question answering and knowledge-grounded dialogue.
+Utilizing LLM’s inherent capabilities for consistency checking is another viable approach.
+SelfCheckGPT (Manakul et al., 2023b) is motivated by the simple idea that LLM tends to have self-consistency if it possesses certain knowledge.
+It uses the scheme of black-box zero-resource hallucination detection and compares multiple sampled responses then measuring consistency.
+Rather than self-consistency, ${}\mathbf{SAC^{3}}$ (Zhang et al., 2023a) performs cross-checking hallucination detection based on sampling.
+By leveraging semantically equivalent question perturbation and cross-model response consistency checking, it can effectively identify the consistency and factuality of the answers.
 KoLA (Yu et al., 2023) use both Wikipedia and continuously collected emerging news and novels to evaluate the world knowledge of LLMs.
 It develops a self-contrast metric by contrasting two completions generated by the same model and gets the similarity score.
-RealTimeQA (Kasai et al., 2024) designs an annotation framework and submission workflow to support the updated evaluation system.
-It tests on both open- and closed-book QA systems and shows the importance of QA systems to be keep-dated.
-ERBench (Oh et al., 2024) verified facts through entity-relationship (ER) model.
-It can construct questions using templates from updated ER databases and evaluates LLMs with binary or multi-choice QA tasks.
-Long Context/Generation:
-Despite recent advancements in LLMs enabling them to better handle long texts, hallucination evaluation in long context or generation remains a challenge.
-FACTSCORE (Min et al., 2023) focuses on long-form text generated by LLMs.
-It breaks the generated text into a series of fine-grained atomic facts by InstructGPT and then each atomic fact is assigned a binary label.
-The data is about biographies of people sampled from English Wikipedia.
-Rather than considering whether the fact is globally truthful, FACTSCORE refers to a given source of knowledge.
-In this way, it can avoid information conflictions between different sources.
-BAMBOO (Dong et al., 2023) provides a multi-task long context benchmark for LLMs which also can detect hallucination.
-It uses an entailment model on two synthetic datasets named SenHallu and AbsHallu to check the hallucination level of LLMs.
 RV (Yang et al., 2023) performs zero-resource hallucination detection by reverse validation.
 It focuses on passage-level detection rather than sentence-level, which holds greater practical significance.
-This method first constructs a question for the entity of user concern from LLM’s output, then uses it to query the LLM and compare the answers.
-This paper also publishes a benchmark PHD for the evaluation of passage-level detection methods.
-Factual Reasoning:
 To evaluate the factual reasoning abilities of LLMs, Laban et al.
@@
 This work provides a three-step protocol for inconsistency detection benchmark creation and implements it in a 10-domain benchmark.
-It shows current LLMs are not yet proficient at complex factual reasoning
-Domain Specific:
-Hallucinations generated by LLMs in the medical field can have serious consequences in healthcare applications.
-Therefore MedHalt (Pal et al., 2023) designs a pipeline to implement three fact-related tests: False Confidence Test (FCT), None of the Above Test (Nota), and Fake Questions Test (FQT).
-Through these tests, MedHalt can effectively identify and filter hallucinations within the medical domain.
-Chinese Language:
-Numerous Chinese LLMs have also emerged along with the trend and hallucination is also a crucial problem.
-ChineseFactEval (Wang et al., 2023a) conducts the hallucination evaluation in seven scenarios, covering areas where generative AI may be applied in daily life, including general knowledge, scientific, medical, legal, financial, mathematics, and modern Chinese history scenarios.
-UHGEval (Liang et al., 2023) provides a Chinese-language dataset for LLM hallucination evaluation, generated by Chinese LLMs in news domain.
-The evaluation method uses LLMs for discriminative evaluation (binary classification), selective evaluation (multiple-choice question), and generative evaluation (use existing lexical and semantic metrics).
-HalluQA (Cheng et al., 2023) is a benchmark categorizing hallucinations into imitative falsehoods and factual errors.
-Inspired by TruthfulQA, it summarizes the question patterns and combines them with Chinese culture to create a new dataset.
-It also instructs GPT-4 to act as a judger to get the consistency rate.
-The result shows this benchmark maintains a challenging benchmark while 18 out of 24 LLMs rate below 50%.
-It should be noted that there is a small portion of overlap between Sec.
-4.2 and Sec.
-4.1, meaning that some approaches utilize LLMs to evaluate LLMs (Chern et al., 2023; Zhao et al., 2024) as well.
-For instance, due to its outstanding performance, GPT-4 can do annotation instead of human sometimes (Cheng et al., 2023).
-In contrast to evaluators before LLMs that primarily focus on SF, methods outlined in this section not only can address SF evaluation but also pay more attention to the WF (general fact) aspect.
-This is because LLMs enable open-domain QA and can even leverage search engines for retrieval.
-Thus, assessing fresh facts is also an important aspect.
-Though the latest LLMs can support long-text processing, challenges still persist in analyzing long contexts or generating long outputs, leading to hallucinations caused by long-distance discourse.
-Factual errors can also accumulate in reasoning or multi-hop question answering.
-Furthermore, in other application domains such as the medical field and non-English language, some preliminary explorations have been conducted, which call for more work participation.
+It shows current LLMs are not yet proficient at complex factual reasoning.
+These methods employ LLMs as tools to assess the hallucination level of models, including the LLMs themselves.
+While using multi-step approaches for evaluation offers strong interpretability, it is prone to error accumulation throughout extraction, reasoning, or multi-hop question answering.
+Similarly, having the model generate multiple answers or using multiple models to generate answers relies heavily on the models’ confidence and robustness.
+This dependency can lead to the oversight of high-confidence errors.
+We additionally present four case examples in Appendix D, covering both SF errors and WF errors.
+Table 2 shows that perfect evaluation for different types of hallucination errors remains challenging, and the evaluators also possess varying degrees of transferability.
 5 Conclusion and Future Direction
-Evaluating hallucination in NLG is crucial as it determines the direction and trend for developing a better model.
-Historically, most metrics are task-based, as defining the criteria for task performance is relatively straightforward.
-As NLP research gradually shifts to LLMs, researchers have proposed new demands and challenges.
+Evaluating hallucination in NLG is crucial as it determines the direction and trend for developing better models.
+Historically, most metrics have been task-based, as defining the criteria for task performance is relatively straightforward.
+However, as NLP research gradually shifts to LLMs, researchers have proposed new demands and challenges.
 We summarize several aspects of hallucination evaluation that urgently need to be addressed in the future.
-Comprehensive evaluation Traditional evaluation methods based on a specific task are no longer sufficient to LLMs.
-Researchers have already proposed many comprehensive evaluation dimensions for the general capabilities, such as common sense and reasoning.
-A comprehensive and unified benchmark specialized for hallucination in the future will greatly promote model iteration to address hallucination issues.
-Differentiation between hallucination and error Hallucinations and text errors can be confusing sometimes.
+Hallucination and text error It can be challenging to distinguish between hallucinations and text errors.
 According to the traditional definition of hallucination (smooth but incorrect), any response from a large model that differs from the ground truth can be considered as hallucination, which is obviously unreasonable and can mislead researchers.
-Some works in NMT have already made some progress (Dale et al., 2023), and it’s better for future evaluation methods to accurately identify real hallucinations.
-Interpretabity Identifying fact granularity and analyzing hallucination reasons can provide significant assistance in solving hallucination problems.
-Some reasoning methods have the potential to analyze the underlying causes of hallucinations and have better evaluation.
-Other observing aspects lie in the internal state of model generation (Su et al., 2024), model structures, and even data distributions, which provide more analytical perspectives.
+Some works in NMT have already made progress in this area (Dale et al., 2023), and future evaluation methods should aim to accurately identify real hallucinations.
+Hallucination and imagination Is hallucination always bad?
+Not necessarily.
+In certain contexts, such as chatting about a sci-fi novel, the content is rich with imagination, and the dialogue should be creative as well.
+In this circumstance, the distinction between hallucination and imagination is quite subtle.
+Distinguishing between these two phenomena can help models evaluate more diverse texts.
+Supporting theories So far, the evaluation and detection of hallucination have been limited to observing the final output of the model – detecting hallucination manifested in the text.
+If there is theoretical support for researchers to detect hallucination from internal hidden states, model structures, and even data distributions, it can greatly reduce the cost of model iteration.
+Interpretability Identifying fact granularity and analyzing the reasons behind hallucination can provide significant assistance in solving hallucination problems.
+Some reasoning methods have the potential to analyze the underlying causes of hallucinations and offer better evaluation.
+Other observing aspects lie in the internal state of model generation (Su et al., 2024), which provide more analytical perspectives.
 Long context/generation Current exploration of LLMs in long-context scenarios mainly focuses on long inputs, such as using a complete contract or multi-turn dialogue history as input.
 It is crucial to address hallucinations caused by the model’s difficulty in understanding long inputs.
-However, on the other hand, hallucinations of the type caused by contradictions between the beginning and end of long outputs are also worth exploring (Wei et al., 2024), such as detecting whether a model-generated novel contains inconsistencies in character behavior.
+However, hallucinations caused by contradictions between the beginning and end of long outputs are also worth exploring (Wei et al., 2024), such as detecting whether a model-generated novel contains inconsistencies in character behavior.
 More application views LLMs have more challenging hallucinations in certain scenarios.
-In the multi-lingual scenario, low-resource languages other than English may be more prone to hallucinations, and cross-lingual conversation can face misunderstanding hallucinations due to different cultural backgrounds.
-Moreover, the latest research focuses on expanding LLMs to areas such as autonomous agents, multimodality, and real-world applications, which bring about new types of hallucinations, such as code hallucination (Qian et al., 2023), tool hallucination, visual hallucination, and so on.
-Detecting such hallucinations is very interesting and worthwhile direction to explore.
+In the multilingual scenario, low-resource languages other than English may be more prone to hallucinations, and cross-lingual conversation can face misunderstanding hallucinations due to different cultural backgrounds.
+Moreover, the latest research focuses on expanding LLMs to areas such as autonomous agents, multimodality, and real-world applications, which bring about new types of hallucinations, such as code hallucination (Qian et al., 2023), tool hallucination, visual hallucination, multimodal hallucination, and so on.
+Detecting such hallucinations is a very interesting and worthwhile direction to explore.
+6 Limination
+In this paper, we collect a broad range of related papers and reports, categorized and compared the methods and offered our insights for future directions.
+However, this paper does have some limitations.
+First of all, we did not do comprehensive experiments to revisit the above evaluators, because the evaluators usually focus on different types of hallucinations for various tasks, and it wouldn’t be fair to compare across the categories.
+For example, evaluators for LLMs intend to build their own datasets with human annotation, which vary in categories and schemes.
+Second, we put the content related to fact-checking and human evaluation in Appendix AB.
+We do not have enough pages to provide a thorough explanation of each of them, and detailed descriptions could also blur the focus of this paper.
+At last but not least, the case study we provide in Appendix D only includes a few representative cases on selective models for reference.
+Acknowledgments
 Appendix A Fact-checking
@@
 ✓and ✗ in SF and WF columns mean the aspects that the evaluators focus on.
-Era Name New Dataset Data Source Fact Definition Task Based-model Method Metric SF WF Before LLM with Ref $Fact_{acc}$ WikiFact Wikipedia, Wikidata KB Ent-rel -ent Summ Transformer NER, Binary Relation Classifier P, R, F1 FactCC FactCC CNN/DM, XSumFaith Sent Summ BERT NLI (2-class) Likelihood DAE DAE PARANMT50M Dependency Summ ELECTRA NLI (2-class) Likelihood Maskeval / CNN/DM, WikiLarge, ASSET Word Summ, Simp T5 Word Weighting Weighted Match Score  Guerreiro et al.
-(2023) Haystack WMT2018 DE-EN Text Span NMT Transformer Uncertainty Measure Avg.
+Appendix D Case Study
+Here we present some results of selected evaluators on different kinds for SF or WF errors on summarization data in Table 2.
+The data we used are from XEnt dataset (Cao et al., 2022) and FactCollect (Ribeiro et al., 2022).
+We selected evaluators that use the GPT series and those that do not, covering both models that evaluate SF and WF facets.
+For the models utilizing LLMs, we specifically employed GPT-3.5-turbo.
+Although FacTool is not directly applicable for evaluating summarization tasks, we conducted experiments under its KBQA (Knowledge-Based Question Answering) setup to see its transfer ability.
+The results of different models on these cases show considerable variation.
+In the SF-WF case, only FacTool made an incorrect judgment, which might be attributed to its insufficient transfer ability.
+SelfCheckGPT uses a zero-shot approach in its prompt to assess the consistency, whereas HaluEval’s prompt provides examples for judgment.
+However, the SFE cases indicate that the results of these two evaluators remain unstable.
+For the WFE cases, FacTool provides the correct answers, and surprisingly, WeCheck also made correct judgments.
+Currently, to our best knowledge, no such labeled data is available for full evaluation.
+More accurate data is needed for further experiments to validate the preferences of different evaluators.
+Era Name New Dataset Data Source Fact Definition Task Based-model Method Metric SF WF Before LLM Ref -based $Fact_{acc}$ WikiFact Wikipedia, Wikidata KB Triplet Summ Transformer NER, Classification P, R, F1 FactCC FactCC CNN/DM, XSumFaith Sent Summ BERT NLI (2-class) Likelihood DAE DAE PARANMT50M Dependency Summ ELECTRA NLI (2-class) Likelihood Maskeval / CNN/DM, WikiLarge, ASSET Word Summ, Simp T5 Word Weighting Weighted Match Score  Guerreiro et al.
+(2023) Haystack WMT2018, DE-EN Text Span NMT Transformer Uncertainty Measure Avg.
 Similarity  Dale et al.
-(2023) / Haystack Text Span NMT Transformer Source Contribution Percentage Before LLM Ref Free FEQA FEQA CNN/DM, XSum Sent Span Summ BART (QG), BERT (QA) QG-QA Avg.
+(2023) / Haystack Text Span NMT Transformer Source Contribution Percentage Before LLM Ref -free FEQA FEQA CNN/DM, XSum Sent Span Summ BART (QG), BERT (QA) QG-QA Avg.
 F1 QAGS QAGS CNN/DM, XSum Ent, Noun Phrase Summ BART (QG), BERT (QA) QG-QA Avg.
-Similarity QuestEval / CNN/DM, Xsum Ent, Noun Summ T5 (QG, QA) QG-QA P, R, F1 QAFactEval / SummaC NP Chunk Summ BART (QG), ELECTRA (QA) QG-QA, NLI LERC MQAG / QAGS, XSumFaith, Podcast Assessment, SummEval Sent Span Summ T5 (QG), Longformer (QA) Multi-Choice QA Choice Statistical Distance CoCo / QAGS, SummEval Token, Span, Sent, Doc Summarization BART Counterfactual Estimation Avg.
-Likelihood Diff ExtEval ExtEval CNN/DM Discourse, Coreference, Sentiment Summ SpanBERT, RoBERTa Direct Prediction, Statistic Summation of Sub-scores $Q^{2}$ $Q^{2}$ WOW Sent Span Diag T5 (QG); Albert- Xlarge (QA); RoBERTa (NLI) QG-QA, NLI Likelihood AlignScore / 22 datasets from 7 tasks Sent NLI, QA, paraphrasing, Fact Verification, IR, Semantic Similarity, Summ RoberTa 3-way classification Likelihood WeCheck / TRUE Response Summ, Diag, Para, Fact Check DeBERTaV3 Weakly Supervised NLI Likelihood After LLM tool SCALE ScreenEval LLM, Human Sentence Long Diag Flan-T5 NLI Likelihood  Chen et al.
-(2023b) / SummEval, XSumFaith, Goyal21, CLIFF Response Summarization Flan-T5, code-davinci-002, text-davinci-003, ChatGPT, GPT-4 Vanilla/COT/ Sent-by-Sent Prompt Balanced Acc GPTScore / 37 datasets from 4 tasks Various Summ, Diag, NMT, D2T GPT-2, OPT, FLAN, GPT-3 Direct Assessment Direct Score G-Eval / SummEval, Topical-Chat, QAGS Response Summ, Diag GPT-4 COT, Form-filling Weighted Scores Wang et al.
-(2023b) / 5 datasets from 3 tasks Response Summ, D2T, Story Gen ChatGPT Direct Assessment, Rating Direct score ChainPoll RealHall-closed, RealHall-open COVID-QA, DROP, Open Ass prompts, TriviaQA Response Hallu Detect gpt-3.5-turbo Direct Assessment (2-class) Acc After LLM object FacTool / RoSE, FactPrompts, HumanEval, GSM-Hard, Self-instruct Claim, Response Knowledge-based QA, Code Generation, Math Reasoning, Sci-literature Review ChatGPT Claim Extraction, Query Generation, Tool Querying, Evidence Collection, Agreement Verification P, R, F1 UFO / NQ, HotpotQA, TruthfulQA, CNN/DM, Multi-News, MS MARCO Ent Open-domain/ Web Retrieval-based/ Expert-validated/ Retrieval-Augmented QA, News Fact Generation ChatGPT (gpt-3.5-turbo-1106) Fact Unit Extraction, Fact Source Verification, Fact Consistency Discrimination Avg.
-Sub-scores CONNER / NQ, WoW Sentence Open-domain QA, Knowledge-grounded Dialogue NLI-RoBERTa -large, ColBERTv2 3-way NLI Acc SelfCheckGPT SelfCheckGPT WikiBio Response Hallu Detect GPT-3 Direct Assessment, QA, BERTScore, NLI, Ngram AUC-PR $SAC^{3}$ / HotpotQA, NQ-open Response QA Generation gpt-3.5-turbo, Falcon-7b-instruct, Guanaco-33b Cross-checking, QA Pair Consistency AUROC EigenScore / CoQA, SQuAD, TriviaQA Natural Questions Inner State Open-book QA Closed-book QA LLaMA, OPT Semantic Consistency/ Diversity in Dense Embedding Space AUROC, PCC TruthfulQA TruthfulQA LLM, Human Response Multi-Choice QA, Generation GPT-3-175B Answer Match Score HaluEval Task-specific, General Alpaca, Task datasets ChatGPT Response QA, Summ, Knowledge- grounded Diag, Generation ChatGPT Direct Assessment Acc FACTOR Wiki-/News-/ Expert- FACTOR Wikipedia, Refin- edWeb, ExpertQA Sent Span Generation / FRANK Error Classification likelihood FELM FELM TruthfulQA, Quora, MMLU, GSM8K, ChatGPT, Human Text Span, Claim World Knowledge, Sci and Tech, Math, Writing and Recom- mendation, Reasoning Vicuna, ChatGPT, GPT4 Direct Assessment F1, Balanced Acc FreshQA Never/Slow Fast-changing, false-premise Human Response Generation / Answer Match Acc KoLA KoLA Wikipedia, Updated News and Novels Response Knowledge Memorization /Understanding/Applying /Creating / Self-contrast Answer Match Similarity RealTimeQA RealTimeQA CNN, THE WEEK, USA Today Response Multi-Choice QA, Generation GPT-3, T5 Answer Match Acc, EM, F1 ERBench ERBench Database 5 datasets from Kaggle Ent-Rel Binary/ Multiple -choice QA / Direct Assessment, String Matching Ans/Rat/ Ans-Rat Acc, Hallu Rate FactScore / Biographies in Wikipedia Atomic Fact Generation InstructGPT, ChatGPT, PerplexityAI Binary Classification P BAMBOO SenHallu, AbsHallu 10 datasets from 5 tasks Response Multi-choice tasks, Select tasks ChatGPT Answer Match P, R, F1 RV PHD Human Annotator Ent Generation ChatGPT Construct Query, Access Databases, Entity-Answer Match P, R, F1 SummEdits SummEdits 9 datasets from Summ task Span Summ, Reasoning gpt-3.5-turbo Seed summary verify, Summary edits, Annotation Balanced Acc MedHalt MedHalt MedMCQA, Medqa USMILE, Medqa (Taiwan), Headqa, PubMed Response Reasoning Hallu Test, Memory Hallu Test ChatGPT Answer Match Pointwise Score, Acc ChineseFactEval ChineseFactEval / Response Generation / FacTool, Human annotator Direct Score UHGEval UHGEval Chinese News Websites Keywords Generative/ Discriminative/ Selective Evaluator GPT-4 Answer Match, Similarity Acc, Similarity Score HalluQA HalluQA Human Response Generation GLM-130B, ChatGPT, GPT-4 Direct Assessment Non-hallu Rate
+Similarity QuestEval / CNN/DM, Xsum Ent, Noun Summ T5 (QG, QA) QG-QA P, R, F1 QAFactEval / SummaC NP Chunk Summ BART (QG), ELECTRA (QA) QG-QA, NLI LERC MQAG / QAGS, XSumFaith, Podcast, Assessment, SummEval Sent Span Summ T5 (QG), Longformer (QA) Multi-Choice QA Choice Statistical Distance CoCo / QAGS, SummEval Token, Span, Sent, Doc Summ BART Counterfactual Estimation Avg.
+Likelihood Diff FactGraph FactCollect CNN/DM, XSum Dependency Summ ELECTRA Classification BACC, F1 FactKB FactKB CNN/DM, XSum Triplet Summ RoBERTa Classification BACC, F1 ExtEval ExtEval CNN/DM Discourse, Coreference, Sentiment Summ SpanBERT, RoBERTa Direct Prediction, Statistic Summation of Sub-scores $Q^{2}$ $Q^{2}$ WOW Sent Span Diag T5 (QG), Albert-Xlarge (QA), RoBERTa (NLI) QG-QA, NLI Likelihood FactPush / TRUE Span Diag, Summ, Paraphrase DeBERTa NLI AUC AlignScore / 22 datasets from 7 tasks Sent NLI, QA, Paraphrase, Fact Verification, IR, Semantic Similarity, Summ RobERTa 3-way Classification Likelihood WeCheck / TRUE Response Summ, Diag, Para, Fact Check DeBERTaV3 Weakly Supervised NLI Likelihood After LLM Direct SCALE ScreenEval LLM, Human Sentence Long Diag Flan-T5 NLI Likelihood  Chen et al.
+(2023b) / SummEval, XSumFaith, Goyal21, CLIFF Response Summ Flan-T5, code-davinci-002, text-davinci-003, ChatGPT, GPT-4 Vanilla/COT/ Sent-by-Sent Prompt Balanced Acc GPTScore / 37 datasets from 4 tasks Various Summ, Diag, NMT, D2T GPT-2, OPT, FLAN, GPT-3 Direct Assessment Direct Score G-Eval / SummEval, Topical-Chat, QAGS Response Summ, Diag GPT-4 COT, Form-filling Weighted Scores Wang et al.
+(2023b) / 5 datasets from 3 tasks Response Summ, D2T, Story Gen ChatGPT Direct Assessment, Rating Direct score ChainPoll RealHall-closed, RealHall-open COVID-QA, DROP, Open Ass prompts, TriviaQA Response Hallu Detect gpt-3.5-turbo Direct Assessment (2-class) Acc EigenScore / CoQA, SQuAD, TriviaQA Natural Questions Inner State Open-book QA Closed-book QA LLaMA, OPT Semantic Consistency/ Diversity in Dense Embedding Space AUROC, PCC TruthfulQA TruthfulQA LLM, Human Response Multi-Choice QA, Generation GPT-3-175B Answer Match Score HaluEval Task-specific, General Alpaca, Task datasets ChatGPT Response QA, Summ, Knowledge- grounded Diag, Generation ChatGPT Direct Assessment Acc FACTOR Wiki-/News-/ Expert- FACTOR Wikipedia, Refin- edWeb, ExpertQA Sent Span Generation / FRANK Error Classification likelihood FELM FELM TruthfulQA, Quora, MMLU, GSM8K, ChatGPT, Human Text Span, Claim World Knowledge, Sci and Tech, Math, Writing and Recom- mendation, Reasoning Vicuna, ChatGPT, GPT4 Direct Assessment F1, Balanced Acc FreshQA Never/Slow Fast-changing, false-premise Human Response Generation / Answer Match Acc RealTimeQA RealTimeQA CNN, THE WEEK, USA Today Response Multi-Choice QA, Generation GPT-3, T5 Answer Match Acc, EM, F1 ERBench ERBench Database 5 datasets from Kaggle Ent-Rel Binary/ Multiple -choice QA / Direct Assessment, String Matching Ans/Rat/ Ans-Rat Acc, Hallu Rate FactScore / Biographies in Wikipedia Atomic Fact Generation InstructGPT, ChatGPT, PerplexityAI Binary Classification P BAMBOO SenHallu, AbsHallu 10 datasets from 5 tasks Response Multi-choice tasks, Select tasks ChatGPT Answer Match P, R, F1 MedHalt MedHalt MedMCQA, Medqa USMILE, Medqa (Taiwan), Headqa, PubMed Response Reasoning Hallu Test, Memory Hallu Test ChatGPT Answer Match Pointwise Score, Acc ChineseFactEval ChineseFactEval / Response Generation / FacTool, Human annotator Direct Score UHGEval UHGEval Chinese News Websites Keywords Generative/ Discriminative/ Selective Evaluator GPT-4 Answer Match, Similarity Acc, Similarity Score HalluQA HalluQA Human Response Generation GLM-130B, ChatGPT, GPT-4 Direct Assessment Non-hallu Rate After LLM Pipeline FacTool / RoSE, FactPrompts, HumanEval, GSM-Hard, Self-instruct Claim, Response Knowledge-based QA, Code Generation, Math Reasoning, Sci-literature Review ChatGPT Claim Extraction, Query Generation, Tool Querying, Evidence Collection, Agreement Verification P, R, F1 UFO / NQ, HotpotQA, TruthfulQA, CNN/DM, Multi-News, MS MARCO Ent Open-domain/ Web Retrieval-based/ Expert-validated/ Retrieval-Augmented QA, News Fact Generation ChatGPT (gpt-3.5-turbo-1106) Fact Unit Extraction, Fact Source Verification, Fact Consistency Discrimination Avg.
+Sub-scores CONNER / NQ, WoW Sentence Open-domain QA, Knowledge-grounded Dialogue NLI-RoBERTa -large, ColBERTv2 3-way NLI Acc SelfCheckGPT SelfCheckGPT WikiBio Response Hallu Detect GPT-3 NLI, Ngram, QA, BERTScore, Prompt AUC-PR $SAC^{3}$ / HotpotQA, NQ-open Response QA Generation gpt-3.5-turbo, Falcon-7b-instruct, Guanaco-33b Cross-checking, QA Pair Consistency AUROC KoLA KoLA Wikipedia, Updated News and Novels Response Knowledge Memorization /Understanding/Applying /Creating / Self-contrast Answer Match Similarity RV PHD Human Annotator Ent Generation ChatGPT Construct Query, Access Databases, Entity-Answer Match P, R, F1 SummEdits SummEdits 9 datasets from Summ task Span Summ, Reasoning gpt-3.5-turbo Seed summary verify, Summary edits, Annotation Balanced Acc
 Table 1: Evaluators and Benchmarks meta information.
+Document Summary Note WeCheck SelfCheckGPT HaluEval FacTool SF-WF … Harry Kane has been given the nod by Youssouf Mulumbu for this season’s players’ Player of the Year award The West Brom midfielder has picked Chelsea wideman Eden Hazard for the young player of the year prize Congo international Mulumbu posted his votes for this year’s PFA awards to Twitter on Wednesday Mulumbu challenges QPR defender Yun Suk-Young during West Brom’s 4-1 defeat at The Hawthorns Goalkeepe … The DR Congo international has picked Chelsea wideman Eden Hazard for the young player of the year prize .
+The summary is correct.
+TRUE TRUE TRUE FALSE SF-WFE … Since the end of March, the Vikings’ only wins have been in the Challenge Cup against lower-league sides.
+"We’ve got the personnel and we’ve got the people to spark us back into life," Chris Betts told BBC Radio Merseyside.
+"When we get rolling again I’m sure, or I’m positive, that we can really turn this year around for ourselves." … "The players are hurting and we’ve got to win," added England assistant coach Betts.
+… Widnes Vikings can turn their poor start to the Super League season around if they can find a winning streak, says assistant coach Chris Betts.
+"Chris Betts" is in the document but is incorrect essentially.
+FALSE TRUE TRUE FALSE SFE-WF The panther chameleon was found on Monday by a dog walker in the wooded area at Marl Park.
+It had to be put down after X-rays showed all of its legs were broken and it had a deformed spine.
+RSPCA Cymru said it was an "extremely sad example of an abandoned and neglected exotic pet".
+…… A chameleon has been put down by RSPCA Cymru after it was found injured and abandoned in a Cardiff park.
+The Marl Park is in Cardiff but not mentioned in the document.
+TRUE FALSE TRUE TRUE SFE-WFE A number of men, two of them believed to have been carrying guns, forced their way into the property at Oakfield Drive shortly after 20:00 GMT on Saturday.
+They demanded money before assaulting a man aged in his 50s.
+… Alliance East Antrim MLA Stewart Dickson has condemned the attack.
+… A man has been assaulted by a gang of armed men during a robbery at a house in Ballymena, County Antrim.
+"Ballymena" is neither in the document nor correct according to external knowledge.
+FALSE TRUE TRUE FALSE
+Table 2: Examples of the results from selected evaluators on the SF errors (SFE) and WF errors (WFE).